### v1

# Q-Mamba: On First Exploration of Vision Mamba for Image Quality Assessment

Abstract
In this work, we take the first exploration of the recently popular foundation model, i.e., State Space Model/Mamba, in image quality assessment, aiming at observing and excavating the perception potential in vision Mamba. A series of works on Mamba has shown its significant potential in various fields, e.g., segmentation and classification. However, the perception capability of Mamba has been under-explored. Consequently, we propose Q-Mamba by revisiting and adapting the Mamba model for three crucial IQA tasks, i.e., task-specific, universal, and transferable IQA, which reveals that the Mamba model has obvious advantages compared with existing foundational models, e.g., Swin Transformer, ViT, and CNNs, in terms of perception and computational cost for IQA. To increase the transferability of Q-Mamba, we propose the StylePrompt tuning paradigm, where the basic lightweight mean and variance prompts are injected to assist the task-adaptive transfer learning of pre-trained Q-Mamba for different downstream IQA tasks. Compared with existing prompt tuning strategies, our proposed StylePrompt enables better perception transfer capability with less computational cost. Extensive experiments on multiple synthetic, authentic IQA datasets, and cross IQA datasets have demonstrated the effectiveness of our proposed Q-Mamba.
footnotetext: ${}^{\textrm{\Letter}}$ Corresponding authors.
1 Introduction
Image Quality Assessment (IQA) aims to measure the subjective quality of images aligned with human perception, which has been applied in various visual fields, including visual acquisition, transmission, AIGC [25, 61, 63, 71], and UGC creation [58, 31, 42], etc. Establishing a great IQA metric is necessary to provide the right optimization direction tailored for image processing techniques, i.e., compression [30, 64, 8, 35], enhancement [27, 28, 24, 7], and ensure the perceptual quality of images. Early works on IQA have been achieved by leveraging the natural scene statistics in a hand-crafted manner [44, 46, 60, 52]. With the advancements of deep neural networks (DNNs), learning-based IQA metrics [29, 38, 34, 76, 77, 3] have demonstrated significant potential for low-level perception, which can be roughly categorized into two types based on the pre-traiend backbones: CNN-based and Transformer-based methods.
Although the impressive progress, learning-based IQA are susceptible to inherent limitations of existing pre-trained backbones: (i) the CNNs are skilled at learning local translation-invariant features from images while lacking enough long-range dependency modeling capability, hindering the global quality perception. (ii) The emergence of Vision Transformers presents a great solution to model long-range dependency effectively by leveraging attention mechanisms. However, the quadratic complexity of self-attention operations poses unaffordable computational costs, especially for large-scale image quality assessment. Recently, an innovative foundation model, the State Space Model, particularly its implementation, i.e., Mamba [12] has shown considerable potential in various fields for balancing the computational costs and performances, e.g., segmentation[65, 43] and classification[54, 80]. This raises one interesting question, i.e., “whether Mamba can surpass existing backbones on low-level visual perception", which is still under-explored.
To answer this question, in this work, we initiate a new exploration of the Mamba model within the field of image quality assessment (IQA), and introduce the Q-Mamba, a newly developed IQA framework designed to address three key facets of IQA: task-specific, universal, and transferable image quality assessment. Notably, since the lack of enough IQA dataset, learning-based IQA metrics entail the pre-trained backbone for perception knowledge extraction. Meanwhile, VMamba [39], as the most representative framework, has achieved excellent performance on high-level tasks by employing horizontal and vertical scanning strategies. However, merely excavating the global perception knowledge is not optimal for IQA, since most artifacts affecting the image quality are related to local textures. Consequently, inspired by LocalMamba[18], we have adopted a local window scanning method, which significantly enhances Mamba’s ability to perceive local distortions, thereby demonstrating superior performance. We detail the architecture of Q-Mamba and examine different scanning methods to clarify how the model interacts with and processes image data. Our extensive analysis confirms that Q-Mamba surpasses traditional foundational models, showing superior perceptual accuracy and greater computational efficiency.
Moreover, we have explored the perception transferability of Mamba across different datasets, i.e., different contents, and degradations. We can find that the Mamba-based IQA metric still suffers from severe performance drop when they encounter large domain shifts between synthetic distortions[53, 33], authentic distortions[10, 17], and Artificial Intelligence-Generated Content (AIGC)[25, 61] distortions, etc. To further amplify the transferability of Q-Mamba across a range of IQA applications, we introduce a simple but effective tuning strategy called StylePrompt. This is based on the finding that the domain shifts in IQA tend to correlate with their feature statistics/style [41], such as mean and variance. Concretely, our StylePrompt aims to adaptively adjust the mean and variance of pre-trained Q-Mamba towards the target IQA tasks by setting a group of light-weight learnable $1\times 1\times C$ parameters. Extensive experiments have shown that our StylePrompt greatly facilitates the task-adaptive learning of Q-Mamba, enabling efficient knowledge transfer across diverse IQA tasks with fewer parameter costs.
The main contributions of this paper are summarized as follows:
  • We embark on a novel exploration of the Mamba model within image quality assessment (IQA), and propose the Q-Mamba, a powerful IQA metric for three critical tasks of IQA: task-specific, universal, and transferable image quality assessment. This exploration has demonstrated the superior potential of Mamba for subjective perception, advancing the development of IQA.
  • To improve the perception transferability of the Q-Mamba, we introduce a simple but effective tuning strategy, i.e., StylePrompt. This strategy enable the efficient knowledge transfer of pre-trained Q-Mamba for downstream IQA tasks, while only tuning fewer learnable parameters to adjust the statistics of perception features.
  • Extensive experiments have shown that our Q-Mamba has consistently achieved state-of-the-art results on various prominent IQA datasets compared with existing IQA methods, thereby validating the efficacy of the Mamba model in quality assessment. Moreover, our StylePrompt achieves nearly equivalent performance to full model tuning while utilizing only $4\%$ of the whole parameters, demonstrating its effectiveness in perception knowledge transfer scenarios.
2 Related Work
2.1 Blind Image Quality Assessment (BIQA)
Early BIQA methods relied heavily on manually designed features for quality score regression [44, 46, 60, 52]. However, these handcrafted features were insufficient for addressing the complexity of BIQA tasks. With the advent of deep learning, network architectures capable of powerful feature extraction significantly improved quality assessment tasks, with Convolutional Neural Networks (CNNs) and Vision Transformers being the most prevalent.
CNN-based BIQA. CNNs have demonstrated robust feature extraction capabilities, leading to their widespread adoption in BIQA tasks. Early works like CNNIQA [20] used convolutional models for feature learning and quality regression, substantially outperforming handcrafted features. DBCNN [75] introduced a dual-stream network to address synthetic and authentic distortions separately, integrating these insights for better quality prediction. NIMA [57] and PQR [72] leveraged pre-trained models on ImageNet for quality score prediction, enhancing accuracy through well-established neural architectures. MetaIQA [79] employed meta-learning to adapt to unknown distortions by learning shared priors for various distortion types, while HyperIQA [56] used a hypernetwork to adaptively establish perceptual rules, improving generalization. Despite these advancements, CNNs’ local bias limits their ability to fully exploit both global and local information in BIQA tasks.
Transformer-based BIQA. Transformers offer superior global modeling capabilities compared to CNNs. TReS [11] addressed CNNs’ local bias by capturing local structural information with CNNs and then using Transformers for sequential feature extraction. MUSIQ [21] designed a multi-scale image Transformer architecture capable of handling images with varying sizes and aspect ratios. DEIQT [50] leveraged a Transformer-based BIQA architecture with attention mechanisms to align with human perception, enhancing model performance and reducing prediction uncertainty. However, the quadratic complexity of Transformers presents a challenge, highlighting the need for architectures with linear complexity capable of global modeling for BIQA tasks.
2.2 State Space Models (SSMs)
State-space models, known for their linear complexity in capturing long-range dependencies, have been integrated into deep learning architectures. The Structured State-Space model (S4) [14] was a pioneer in deep state-space modeling for remote dependency modeling. Subsequent advancements [13, 55, 9] further propelled the development of state-space models. Mamba [12], by integrating selection mechanisms and hardware-aware algorithms, has shown effective long-range modeling capabilities with linear complexity growth.
Initially focused on NLP tasks, Mamba has rapidly expanded into other domains. Vim [80] introduced a bidirectional SSM block for visual representation learning, achieving performance comparable to ViT [5]. VMamba [39] introduced a cross-scan module to traverse spatial domains and convert non-causal visual images into ordered block sequences, maintaining linear complexity while retaining global receptive fields. LocalMamba [18] employed a window-based scanning approach to integrate local inductive biases, enhancing the visual Mamba model. These advancements validate Mamba’s efficacy in visual tasks, leading to its application in image classification [39, 80, 47, 67, 1], video understanding [62, 4, 26], image restoration [15, 54, 78], point cloud analysis [32, 74, 37], and biomedical image segmentation [43, 65, 51, 36]. This paper explores Mamba’s potential in visual quality perception tasks, providing a simple yet effective baseline for future quality perception research.
3 Method
3.1 Prelimineries
State Space Model (SSM).
Traditional State Space Models (SSMs) utilize first-order differential equations to establish a linear time-invariant system, translating one-dimensional input sequences $x(t)\in\mathbb{R}$ into outputs through hidden states $h(t)$ and $y(t)$, described by:
$\begin{gathered}h^{\prime}(t)=Ah(t)+Bx(t)\\ y(t)=Ch(t)\end{gathered}$ (1)
For compatibility with deep learning’s sequential input requirements, these models undergo a transition to a discrete framework within the S4 model, using discretization parameters $\Delta$ to modify $A$ and $B$ into $\overline{A}$ and $\overline{B}$:
$\begin{gathered}h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t}\\ y_{t}=Ch_{t}\end{gathered}$ (2)
$\begin{gathered}\overline{A}=\exp(\Delta A)\\ \overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B\end{gathered}$ (3)
Mamba[12], evolving from these foundations, introduces a selection algorithm for dynamic adaptation of $A$, $B$, and $C$ based on input, enhancing model response to complex sequences and contextual variations. However, this necessitates abandoning traditional convolutional methods for kernel fusion and parallel scanning to maintain computational efficiency, effectively scaling with sequence length and providing advanced capabilities in sequence generation for language models.
Figure 1: Overview of the Q-Mamba architecture and description of the StylePrompt.
3.2 Exploring Mamba for Perception
3.2.1 Overall Framework
The overall architecture of our Q-Mamba model is illustrated in Figure1. To successfully incorporate the Mamba model into visual perception tasks, we implemented the Q-Mamba architecture following the construction method of the VMamba’s visual Mamba blocks. This architecture consists of multiple layers of VMamba blocks, with the SS2D block serving as the core module, which facilitates information transmission through a residual network. The residual network incorporates convolutional layers for feature extraction and activation layers for computing gating signals. We employed multiple network stages, each comprising a downsampling layer and a VMamba Block, which construct multi-level representations at varying resolutions through downsampling, thus extracting richer perceptual features. To explore the impact of model size on quality perception tasks, we experimented with three different scales of the model: QMamba-Tiny, QMamba-Small, and QMamba-Base.
3.2.2 Perception with Local Scanning
Originally, Mamba demonstrated exceptional proficiency in NLP tasks, characterized by inherently causal sequential inputs. However, when adapted to visual tasks, which lack corresponding causal frameworks, the model encountered substantial difficulties. It struggled to effectively discern the complex interdependencies among image pixels, thereby impeding its learning processes.
VMamba[39] employs a novel bidirectional horizontal and vertical scanning mechanism, illustrated in Figure 2(a), to convert two-dimensional data into one-dimensional inputs for model ingestion. Although this method facilitates global modeling of pixel sequences, it unfortunately causes locally adjacent tokens to be spaced significantly apart, thus failing to capture essential local dependencies. This limitation becomes particularly acute in quality assessment tasks, where accurate distortion detection is dependent on the granularity of local detail.
Drawing inspiration from LocalMamba[18], we have implemented a window-based scanning strategy, as depicted in Figure 2(b), to enhance local perceptual accuracy. This involves conducting horizontal scans within each window followed by comprehensive window-level horizontal scans, with a parallel approach adopted in the vertical dimension. This method markedly enhances VMamba’s capability to detect local quality details. Despite the model’s input sequence undergoing downsampling and aggregation, this scanning approach retains an ability for global perception, ensuring a robust analytical balance that effectively captures both detailed local nuances and broader contextual information.
Figure 2: Scanning Methodology Illustration. (a) VMamba’s[39] method flattens 2D data into 1D, impairing connectivity by distancing adjacent tokens. (b) Our local scanning method[18] scans within and across windows, placing semantically similar and distortion-related tokens closer, as shown in the blue boxes.
3.3 Tuning the Mamba with StylePrompt
Although the Mamba architecture offers reduced computational complexity compared to other models, achieving higher performance still necessitates a substantial number of parameters, which poses a challenge for efficient and transferable learning critical in Image Quality Assessment (IQA) tasks. Prompt tuning[19, 23], a common method for efficient fine-tuning, presents a promising approach. However, conventional prompt tuning techniques involving convolutional or attention operations still require significant parameter overhead[49]. Considering the need for transfer learning in IQA tasks to adapt specifically to the current distortion style, we propose the StylePrompt to enhance the transferability of IQA capabilities. Our method achieves results comparable to full-parameter fine-tuning with a minimal parameter increase.We illustrate the StylePrompt we used in Figure1. Our StylePrompt consists of two parts, which are described in detail below:
3.3.1 StylePrompt Generation (SPG)
In a multi-stage network architecture, as images progress through each stage of the network, we learn a set of prompts $P_{s}\in\mathbb{R}^{N\times 1\times 1\times{C}}$, containing N prompt components designed to generate affine parameters for style adaptation. These prompts are specifically utilized to inject the distortion style of the current image into the features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, facilitating learning of the style pertinent to the current domain. We have devised a straightforward StylePrompt Generation stage where the prompt components $P_{s}$ are initially formed as a set of learnable parameters, interacting with the incoming features to embed distortion information.
To enable the prompt components to extract specific style information from the input stream dynamically, we predict the weights for different prompt components based on the input features. This process involves performing a global pooling on the current layer’s features followed by applying a softmax function to obtain the weights for the prompt group. These weights are then applied to the multiple prompt components to amalgamate them into a new prompt $P_{f}$, effectively encapsulating the current style information. The operation can be succinctly summarized by the following formula:
$P_{f}=\sum_{c=1}^{N}w_{s}P_{s},\quad w_{s}=\text{Softmax}(\text{Conv1x1}(\text{GAP}(F_{i})))$ (4)
3.3.2 StylePrompt Injection (SPI)
In the StylePrompt Generation phase, we created the fused prompt $P_{f}$, which contains the distortion style information of the current domain. During the StylePrompt Injection process, this style information is injected into the original features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$ produced by the current layer. We adapt the distribution of the original features to the style by learning affine parameters. Specifically, as each layer utilizes a fixed-size prompt component, the fused Prompt generated in the StylePrompt Generation phase is matched to the channel dimensions of the current features through linear layers designed to align dimensions. This vector is then processed through two specific linear layers whose primary function is to adjust dimensions, thereby producing the affine parameters $\gamma_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$ and $\beta_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$. These parameters are used to inject the distorted style information solely on the channel dimension into the original feature distribution. The process of our StylePrompt Injection can be summarized as follows:
$\gamma_{v}=\text{Linear}_{\gamma}(\text{Conv}(P_{f}))$ (5)
$\beta_{v}=\text{Linear}_{\beta}(\text{Conv}(P_{f}))$ (6)
$F{{}^{\prime}}_{i}=F_{i}\cdot(1+\gamma)+\beta$ (7)
Specifically, $F_{i}^{\prime}$ represents the original features after the injection of style information. These enhanced features will serve as the new input to the subsequent layer of the network.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We conducted foundational experiments on ten popular IQA datasets, which include four synthetic datasets: LIVE[53], CSIQ[22], TID2013[48], and KADID[33]; four authentic datasets: LIVEC[10], KonIQ[17], LIVEFB[68], and SPAQ[6]; and two AIGC datasets: AIGC2023[61] and AGIQA3K[25]. In our single-task experiments, we utilized all ten datasets for testing. We report the results for the eight datasets consisting of synthetic and authentic datasets in the main text, including comparisons with existing SOTA BIQA methods and performance comparisons of Swin Transformer, ViT, and ResNet across different parameter versions. In the appendix, we report the comparative results of the two AIGC datasets against the aforementioned model architectures. In our universal IQA and transferable IQA experiments, we selected six datasets for training and testing, which include two synthetic datasets (LIVE[53], KADID[33]), two authentic datasets (LIVEC[10], KonIQ[17]), and two AIGC datasets (AIGC2023[61], AGIQA3K[25]).
4.1.2 Evaluation Criteria
The evaluation metrics employed in our study are the widely utilized Pearson Linear Correlation Coefficient (PLCC) and Spearman’s Rank Correlation Coefficient (SRCC), both of which range from 0 to 1. Values approaching 1 denote a higher degree of prediction relevance.
4.1.3 Experimental Details
Our experimental methodology closely adheres to the training strategy outlined in DEIQT[50], where input images are randomly cropped into ten patches of 224×224 resolution. We utilized three variants of the VMamba architecture: VMamba-B, VMamba-S, and VMamba-T. Both VMamba-B and VMamba-S feature an encoder with a depth of 15 layers, where VMamba-B incorporates an embedding dimension of 128, and VMamba-S has an embedding dimension of 96. In contrast, VMamba-T is characterized by a reduced depth of 4 layers and an embedding dimension of 96. Training procedures leveraged weights pre-trained on the ImageNet-1K dataset, spanning across a total of 9 epochs, with batch sizes adjusted according to the respective dataset sizes. All experiments were conducted utilizing multiple NVIDIA RTX 4090 GPUs.
Table 1: Performance comparison for task-specific IQA. Bold indicates the top two results.
  Model LIVE CSIQ TID2013 KADID LIVEC KonIQ LIVEFB SPAQ
  FLOPs PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
     ILNIQE[73] - 0.906 0.902 0.865 0.822 0.648 0.521 0.558 0.534 0.508 0.508 0.537 0.523 0.332 0.294 0.712 0.713 0.618
BRISQUE[45] - 0.944 0.929 0.748 0.812 0.571 0.626 0.567 0.528 0.629 0.629 0.685 0.681 0.341 0.303 0.817 0.809 0.664
WaDIQaM[2] - 0.955 0.960 0.844 0.852 0.855 0.835 0.752 0.739 0.671 0.682 0.807 0.804 0.467 0.455 - - 0.763
DBCNN[75] - 0.971 0.968 0.959 0.946 0.865 0.816 0.856 0.851 0.869 0.851 0.884 0.875 0.551 0.545 0.915 0.911 0.852
TIQA[70] - 0.965 0.949 0.838 0.825 0.858 0.846 0.855 0.850 0.861 0.845 0.903 0.892 0.581 0.541 - - 0.829
MetaIQA[79] - 0.959 0.960 0.908 0.899 0.868 0.856 0.775 0.762 0.802 0.835 0.856 0.887 0.507 0.540 - - 0.815
P2P-BM[69] - 0.958 0.959 0.902 0.899 0.856 0.862 0.849 0.840 0.842 0.844 0.885 0.872 0.598 0.526 - - 0.835
HyperIQA[56] - 0.966 0.962 0.942 0.923 0.858 0.840 0.845 0.852 0.882 0.859 0.917 0.906 0.602 0.544 0.915 0.911 0.858
TReS[11] - 0.968 0.969 0.942 0.922 0.883 0.863 0.858 0.859 0.877 0.846 0.928 0.915 0.625 0.554 - - 0.858
MUSIQ[21] - 0.911 0.940 0.893 0.871 0.815 0.773 0.872 0.875 0.746 0.702 0.928 0.916 0.661 0.566 0.921 0.918 0.832
DEIQT[50] - 0.982 0.980 0.963 0.946 0.908 0.892 0.887 0.889 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919 0.884
LoDa[66] - 0.979 0.975 - - 0.901 0.869 0.936 0.931 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925 0.882
     ResNet-50(23.51M) 4.11G 0.879 0.884 0.861 0.841 0.747 0.686 0.784 0.786 0.868 0.831 0.908 0.886 0.313 0.269 0.907 0.907 0.772
ResNet-101(42.50M) 7.83G 0.918 0.921 0.891 0.867 0.779 0.727 0.722 0.719 0.862 0.824 0.918 0.904 0.420 0.347 0.908 0.906 0.790
ResNet-152(58.15M) 11.53G 0.926 0.927 0.923 0.899 0.765 0.717 0.764 0.760 0.859 0.816 0.919 0.898 0.433 0.353 0.907 0.907 0.798
     ViT-T(5.52M) 1.26G 0.786 0.792 0.725 0.717 0.728 0.699 0.832 0.836 0.777 0.730 0.852 0.852 0.521 0.461 0.896 0.896 0.756
ViT-S(21.67M) 4.61G 0.900 0.896 0.832 0.815 0.873 0.859 0.893 0.894 0.831 0.799 0.922 0.905 0.539 0.443 0.919 0.917 0.827
ViT-B(85.80M) 17.58G 0.961 0.955 0.924 0.912 0.904 0.905 0.910 0.908 0.875 0.837 0.913 0.895 0.491 0.452 0.914 0.912 0.854
     Swin-T(27.52M) 4.51G 0.879 0.883 0.865 0.847 0.937 0.925 0.923 0.922 0.880 0.845 0.901 0.881 0.476 0.453 0.922 0.919 0.841
Swin-S(48.84M) 8.77G 0.883 0.896 0.884 0.874 0.931 0.918 0.895 0.894 0.907 0.884 0.931 0.914 0.476 0.433 0.918 0.915 0.847
Swin-B(86.74M) 15.47G 0.945 0.948 0.941 0.935 0.942 0.933 0.934 0.932 0.892 0.858 0.945 0.932 0.507 0.471 0.923 0.921 0.872
     QMamba-T (27.99M) 4.47G 0.959 0.959 0.940 0.918 0.951 0.945 0.934 0.930 0.898 0.866 0.941 0.925 0.675 0.581 0.934 0.929 0.893
QMamba-S (49.37M) 8.71G 0.962 0.965 0.921 0.903 0.957 0.955 0.934 0.933 0.903 0.874 0.943 0.930 0.677 0.573 0.932 0.927 0.893
QMamba-B (87.53M) 15.35G 0.960 0.961 0.908 0.889 0.953 0.949 0.935 0.932 0.908 0.876 0.943 0.930 0.675 0.579 0.933 0.929 0.891
     LQMamba-T(29.87M) 4.44G 0.958 0.959 0.935 0.916 0.952 0.950 0.938 0.923 0.903 0.863 0.943 0.928 0.672 0.574 0.933 0.927 0.892
LQMamba-S(52.91M) 8.66G 0.962 0.964 0.933 0.914 0.955 0.949 0.941 0.928 0.907 0.882 0.946 0.934 0.676 0.574 0.933 0.929 0.895
LQMamba-B(93.79M) 15.30G 0.959 0.951 0.915 0.889 0.965 0.964 0.943 0.941 0.913 0.888 0.947 0.933 0.675 0.582 0.934 0.929 0.896
4.2 A comparison between different IQA backbones
4.2.1 Task-specific IQA
We conducted comprehensive training and testing across the ten datasets previously introduced, drawing analytical comparisons based on results reported by existing methods. Given the predominance of existing BIQA methods targeting synthetic and distorted datasets, we present in Table 1 a detailed comparative analysis of the performance of state-of-the-art (SOTA) methods alongside popular architectures such as ResNet[16], ViT[5], and Swin Transformer[40], against our Q-Mamba architecture. This comparison elucidates performance discrepancies and computational complexities across different parameter configurations. Our findings demonstrate that the LocalQMamba configuration achieves optimal performance across six datasets and, when juxtaposed with other models of comparable parameter scales, it exhibits lower Gflops, thereby confirming the efficacy and reduced computational complexity of the Mamba architecture. Comparative results for the two AIGC datasets are documented in the appendix in Table 7.
4.2.2 Universal IQA
The efficacy of the Q-Mamba architecture was rigorously assessed across a spectrum of datasets, adopting a joint training protocol on six varied datasets, encompassing two synthetic datasets (LIVE, KADID), two authentic datasets (LIVEC, KonIQ), and two AIGC datasets (AIGC2023, AGIQA3K). Subsequent to this joint training phase, we evaluated the performance outcomes, with the findings depicted in Table 2. These results highlight Q-Mamba’s robust multi-tasking capabilities. In comparison to several established foundational models, Q-Mamba consistently delivered notable performance across the majority of the datasets, thereby substantiating its versatility and effectiveness in managing generalized tasks with high efficacy.
Figure 3: t-SNE clustering of different backbones for various types of distortions
4.2.3 Analysis
To elucidate the reasons behind the VMamba architecture’s notable performance on task-specific IQA and universal IQA tasks, we engaged a subset of the KADID dataset featuring seven distinct types of distortions. We conducted t-SNE[59] dimensionality reduction clustering on the features processed through four disparate model architectures, as illustrated in Figure 3. The clustering outcomes demonstrate that the VMamba architecture exhibits exceptional distortion perception capabilities, effectively distinguishing various types of distortions and surpassing the performance of the ViT structure. Conversely, the ResNet and Swin Transformer architectures did not manifest significant abilities in aggregating different distortion types. Consequently, it becomes apparent that VMamba’s exemplary performance across diverse quality assessment tasks can be attributed, in part, to its robust distortion perception capabilities.
Table 2: Performance comparison for universal IQA. Bold indicates the top two results.
  Train LIVE & KADID & LIVEC & KonIQ & AGIQA3K & AIGCIQA2023
  Test LIVE KADID LIVEC KonIQ AGIQA3K AIGCIQA2023
  GFLOPS PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
     ResNet-50 4.11G 0.916 0.908 0.868 0.869 0.885 0.841 0.896 0.869 0.884 0.813 0.821 0.815 0.865
ResNet-101 7.83G 0.927 0.921 0.870 0.863 0.885 0.848 0.906 0.878 0.897 0.841 0.843 0.833 0.876
ResNet-152 11.53G 0.936 0.932 0.872 0.866 0.890 0.853 0.911 0.889 0.901 0.841 0.852 0.844 0.882
ViT-T 1.26G 0.878 0.878 0.829 0.828 0.811 0.773 0.824 0.805 0.824 0.745 0.767 0.766 0.811
ViT-S 4.61G 0.921 0.918 0.920 0.917 0.883 0.828 0.911 0.892 0.885 0.822 0.825 0.824 0.879
ViT-B 17.58G 0.941 0.935 0.903 0.901 0.884 0.844 0.912 0.899 0.890 0.826 0.848 0.832 0.885
Swin-T 4.51G 0.927 0.927 0.925 0.921 0.889 0.863 0.929 0.918 0.897 0.840 0.834 0.828 0.892
Swin-S 8.77G 0.946 0.949 0.920 0.916 0.913 0.884 0.929 0.912 0.902 0.844 0.861 0.852 0.902
Swin-B 15.47G 0.945 0.942 0.940 0.937 0.916 0.886 0.940 0.928 0.905 0.847 0.880 0.865 0.911
QMamba-T 4.47G 0.923 0.923 0.938 0.932 0.898 0.863 0.932 0.917 0.906 0.851 0.835 0.830 0.896
QMamba-S 8.71G 0.938 0.941 0.944 0.938 0.905 0.875 0.937 0.923 0.909 0.855 0.853 0.841 0.905
QMamba-B 15.35G 0.922 0.924 0.947 0.942 0.911 0.883 0.939 0.924 0.908 0.857 0.852 0.840 0.904
LQMamba-T 4.44G 0.929 0.926 0.943 0.939 0.899 0.863 0.936 0.921 0.908 0.853 0.840 0.828 0.899
LQMamba-S 8.66G 0.935 0.936 0.942 0.938 0.902 0.870 0.938 0.923 0.912 0.860 0.859 0.843 0.905
LQMamba-B 15.30G 0.941 0.939 0.944 0.941 0.916 0.889 0.939 0.927 0.910 0.859 0.870 0.857 0.911
Table 3: Performance comparison for transferable IQA.
  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K KonIQ LIVEC KADID LIVE
    Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
    Without_tuning 0.535 0.501 0.726 0.744 0.789 0.784 0.726 0.744 0.553 0.549 0.595 0.550 0.600 0.612 0.675 0.649 0.636 0.587 0.667 0.637 0.460 0.425 0.815 0.838 0.642
Lin_Probe(R) 0.606 0.581 0.783 0.811 0.824 0.804 0.769 0.695 0.871 0.848 0.785 0.751 0.803 0.788 0.771 0.723 0.818 0.792 0.790 0.761 0.570 0.540 0.799 0.826 0.755
Full_tuning (93.79M)
0.936 0.930 0.941 0.940 0.885 0.863 0.910 0.851 0.943 0.928 0.899 0.873 0.880 0.859 0.910 0.856 0.943 0.927 0.910 0.878 0.937 0.932 0.938 0.933 0.908
StylePrompt (3.83M)
0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
StylePrompt & R 0.921 0.913 0.944 0.945 0.877 0.853 0.906 0.847 0.931 0.912 0.889 0.860 0.876 0.857 0.905 0.849 0.929 0.909 0.899 0.859 0.918 0.911 0.926 0.927 0.898
4.3 Efficent transfer learning for Mamba-based IQA
In the context of transferable IQA tasks, we engaged in domain-specific training utilizing synthetic datasets (LIVE, KADID), authentic datasets (LIVEC, KonIQ), and AIGC datasets (AIGC2023, AGIQA3K). Each dataset underwent subsequent evaluations in the other two domains. We employed the StylePrompt technique, whereby we maintained the architecture’s integrity during cross-domain transfer by freezing it entirely, and solely fine-tuned the StylePrompt, utilizing merely $4\%$ of the total parameter volume. This approach achieved results that were nearly equivalent to those obtained through full-parameter training. The outcomes, as delineated in Table 3, present a comparative analysis of various fine-tuning strategies, conclusively validating the efficacy and efficiency of our proposed StylePrompt method in facilitating transferable IQA tasks.
4.4 Ablation Studies
4.4.1 Different scanning
Examining the outcomes from task-specific assessment in Table 1 and universal assessment in Table 2 , models equipped with local scanning have shown a definitive advantage over those utilizing cross-scanning techniques. This observation has led to an exploration of the effectiveness of local scanning methods.
4.4.2 Different model size
In our empirical analysis, as documented in Table 1 and Table 2, we discern that while QMamba-Base exhibits exceptionally robust quality perception capabilities, QMamba-Small either matches or exceeds the performance of QMamba-Base across the majority of the datasets. Although QMamba-Tiny displays a modest decline in performance metrics, it still delivers results that are competitive with current state-of-the-art (SOTA) methods, solely utilizing the capabilities of QMamba-Tiny.
Table 4: Ablation study on different prompt tuning strategies
  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K KonIQ LIVEC KADID LIVE
    Style and Prompt method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
    SSF(6.1M) 0.718 0.700 0.735 0.769 0.804 0.801 0.714 0.667 0.863 0.829 0.643 0.598 0.668 0.675 0.714 0.685 0.866 0.828 0.730 0.706 0.724 0.708 0.820 0.856 0.743
Crossattn_Prompt(12.17M) 0.843 0.830 0.898 0.897 0.783 0.752 0.821 0.706 0.843 0.816 0.641 0.612 0.802 0.782 0.833 0.726 0.842 0.818 0.644 0.616 0.831 0.815 0.894 0.897 0.789
Conv_Prompt(28.33M) 0.911 0.910 0.945 0.946 0.889 0.820 0.890 0.825 0.897 0.881 0.797 0.761 0.864 0.827 0.861 0.828 0.898 0.884 0.805 0.756 0.901 0.899 0.932 0.933 0.869
StylePrompt(ours)(3.83M) 0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
4.4.3 Different prompt tuning strategies
To validate the effectiveness of the StylePrompt Generation (SPG) process, we tested the direct learning of a set of affine parameters, $\gamma$ and $\beta$, to modulate the original features. The results in Table 4 demonstrate the outcomes of transfer learning using only the learned affine parameters, substantiating the effectiveness of our StylePrompt Generation phase.
Furthermore, to verify the efficacy of different prompt interaction methods within the StylePrompt, we evaluated various prompt interaction strategies with the original features. These included convolutional prompt interaction and cross-attention prompt interaction. Experimental results indicated that the interaction method employed by the StylePrompt with the original features achieves the best transfer results with minimal parameter usage. This further validates the effectiveness of our StylePrompt Injection (SPI) phase.
Table 5: Ablation study on the number of prompt
  Num of Prompt PLCC_Average SRCC_Average Average
  N=1 0.905 0.881 0.8928
N=2 0.905 0.882 0.8936
N=4 0.907 0.884 0.8951
N=6 0.911 0.890 0.9006
N=8 0.905 0.883 0.8939
N=10 0.906 0.883 0.8942
Table 6: Ablation study on the shape of prompts
  Prompt_shape(H,W) PLCC_Average SRCC_Average Average
  (1,1) 0.911 0.890 0.901
(7,7) 0.909 0.886 0.897
(14,14) 0.909 0.886 0.897
(28,28) 0.907 0.883 0.895
layer_HW 0.908 0.884 0.896
4.4.4 Different Prompt Designs
We investigated the impact of the number of prompt components on performance. Our findings revealed that setting the number of prompts to six yields optimal performance. Furthermore, we explored whether prompts with different spatial dimensions and sizes contribute to enhanced capabilities. Ultimately, we found that a spatial size of (1,1), which solely consists of the channel dimension, yields superior performance. We report the average results of these two ablation studies across multiple datasets in Tables 5 and Table 6. More detailed data will be provided in the appendix Table 8 and Table 9.
5 Conclusion
In this paper, we introduced Q-Mamba, a novel exploration of the Mamba model in the IQA field, which aims to enhance image quality assessment (IQA) from three perspectives: task-specific, universal, and transferable IQA. Our extensive experiments have demonstrated that Q-Mamba has great potential to outperform existing foundational models in IQA, including Swin Transformer, ViT, and CNNs, in both perceptual accuracy and computational efficiency. To improve the perception transferability of Q-Mamba, we developed StylePrompt, a powerful tuning mechanism, to enable the effective perception knowledge transfer of Q-Mamba to various downstream IQA tasks even with large domain shifts, by exploiting only extra $4\%$ of whole parameters to adjust the statistics of perception features of Q-Mamba. We believe these findings can prompt the further development of IQA and pave the way for future research of Mamba-based quality assessment.
Appendix A Appendix / supplemental material
A.1 Additional experimental results
A.1.1 Task-specific results on AIGCIQA
We present the test results of various backbones on the AIGC datasets in Table 7, with bold indicating the best results.
Table 7: Results of AIGC on various backbones
  Model AGIQA3K AIGCIQA2023
  FLOPs PLCC SRCC PLCC SRCC PLCC_Average SRCC_Average Average
     ResNet-50(23.51M) 4.11G 0.901 0.840 0.795 0.797 0.848 0.819 0.833
ResNet-101(42.50M) 7.83G 0.907 0.847 0.834 0.831 0.871 0.839 0.855
ResNet-152(58.15M) 11.53G 0.901 0.832 0.841 0.834 0.871 0.833 0.852
ViT-T(5.52M) 1.26G 0.865 0.787 0.760 0.766 0.813 0.777 0.795
ViT-S(21.67M) 4.61G 0.891 0.819 0.842 0.822 0.867 0.821 0.844
ViT-B(85.80M) 17.58G 0.897 0.830 0.853 0.835 0.875 0.833 0.854
Swin-T(27.52M) 4.51G 0.906 0.847 0.867 0.844 0.887 0.846 0.866
Swin-S(48.84M) 8.77G 0.908 0.849 0.875 0.857 0.892 0.853 0.872
Swin-B(86.74M) 15.47G 0.909 0.852 0.886 0.863 0.898 0.858 0.878
QMamba-T (27.99M) 4.47G 0.913 0.858 0.888 0.873 0.901 0.866 0.883
QMamba-S (49.37M) 8.71G 0.912 0.858 0.889 0.875 0.901 0.867 0.884
QMamba-B (87.53M) 15.35G 0.914 0.861 0.886 0.868 0.900 0.865 0.882
LQMamba-T(29.87M) 4.44G 0.914 0.862 0.884 0.868 0.899 0.865 0.882
LQMamba-S(52.91M) 8.66G 0.913 0.864 0.888 0.869 0.901 0.867 0.884
LQMamba-B(93.79M) 15.30G 0.915 0.858 0.888 0.871 0.902 0.865 0.883
A.1.2 Detailed Results of Ablation Studies
We report the detailed data of our experiments on Prompt design in Table 8 and Table 9.
Table 8: Ablation study results for different numbers of prompts (N).
  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K KonIQ LIVEC KADID LIVE
  PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
     N=1 0.914 0.907 0.943 0.944 0.877 0.850 0.900 0.833 0.924 0.906 0.873 0.846 0.872 0.850 0.898 0.834 0.924 0.907 0.890 0.855 0.904 0.899 0.940 0.937 0.8928
N=2 0.913 0.905 0.938 0.939 0.874 0.851 0.895 0.837 0.930 0.912 0.886 0.853 0.871 0.851 0.892 0.835 0.927 0.905 0.902 0.865 0.911 0.906 0.926 0.922 0.8936
N=4 0.920 0.913 0.944 0.949 0.875 0.851 0.904 0.840 0.929 0.913 0.877 0.849 0.876 0.857 0.904 0.842 0.927 0.904 0.885 0.847 0.913 0.907 0.925 0.931 0.8951
N=6 0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.9006
N=8 0.920 0.914 0.930 0.934 0.874 0.851 0.903 0.841 0.930 0.912 0.873 0.843 0.876 0.854 0.899 0.841 0.927 0.906 0.889 0.859 0.918 0.912 0.922 0.926 0.8939
N=10 0.919 0.913 0.930 0.936 0.877 0.856 0.901 0.840 0.929 0.910 0.880 0.849 0.877 0.853 0.900 0.840 0.925 0.906 0.889 0.851 0.918 0.914 0.923 0.924 0.8942
Table 9: Ablation study for different prompt shape
  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K KonIQ LIVEC KADID LIVE
     (H,W) PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
     (1,1) 0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
(7,7) 0.924 0.917 0.947 0.948 0.876 0.852 0.906 0.844 0.931 0.912 0.881 0.848 0.875 0.857 0.902 0.843 0.928 0.906 0.890 0.855 0.921 0.915 0.927 0.931 0.897
(14,14) 0.924 0.917 0.947 0.948 0.876 0.852 0.906 0.844 0.931 0.912 0.881 0.848 0.875 0.857 0.902 0.843 0.928 0.906 0.890 0.855 0.921 0.915 0.927 0.931 0.897
(28,28) 0.917 0.909 0.948 0.947 0.871 0.849 0.905 0.847 0.929 0.909 0.876 0.845 0.873 0.855 0.903 0.841 0.928 0.905 0.895 0.860 0.902 0.898 0.932 0.931 0.895
layer_HW 0.921 0.914 0.948 0.948 0.880 0.854 0.905 0.844 0.929 0.909 0.876 0.850 0.879 0.859 0.903 0.844 0.928 0.906 0.891 0.858 0.913 0.908 0.920 0.919 0.896
A.2 Limitations
Q-Mamba heavily relies on pre-trained weights from ImageNet-1K, which may limit its applicability to domains with significantly different data distributions. Future work could explore pre-training on more diverse datasets to improve generalization capabilities. Despite the efficiency improvements brought by the StylePrompt, the computational demands for training and deploying Q-Mamba on very large-scale datasets or in real-time applications might still be substantial. Investigating methods to further reduce computational complexity without sacrificing performance could be beneficial. The current study focuses on specific types of distortions, and there may be other types of distortions that have not been sufficiently explored. Expanding the evaluation to cover a broader range of distortions could provide a more comprehensive validation of Q-Mamba’s robustness. Given that synthetic datasets might not fully capture the complexity and variability of authentic data, there is a risk of overfitting to these synthetic examples. Further evaluations on more diverse and extensive authentic datasets would help ensure the model’s robustness and practical applicability. While Q-Mamba shows promising results in cross-domain transferability, its effectiveness across vastly different domains (e.g., medical imaging versus natural images) has not been thoroughly validated. Further studies are needed to test and possibly adapt Q-Mamba for such diverse applications.
A.3 Broader Impacts
Improving image quality assessment (IQA) through Q-Mamba allows users across various platforms to enjoy higher quality images, enhancing user experience in social media, streaming services, and other digital platforms, thus leading to greater satisfaction and engagement. The enhancement of image quality can significantly benefit individuals with visual impairments, as better image clarity and quality make content more accessible, thereby improving the inclusivity of digital media. Additionally, the computational efficiency of Q-Mamba can result in more efficient image processing workflows, reducing the time and energy required for image enhancement, compression, and transmission. This can lead to cost savings and environmental benefits due to lower energy consumption. Moreover, the transferability of Q-Mamba across different domains can improve image quality in fields such as medical imaging, satellite imagery, and autonomous driving, thereby enhancing the accuracy and reliability of critical applications.
While Q-Mamba presents these positive societal impacts, it is important to acknowledge and mitigate potential negative effects. Enhancing image quality and clarity may inadvertently lead to privacy concerns, as higher quality images can capture more detailed information about individuals and environments. However, our work adheres to ethical guidelines and uses publicly available, open-source datasets, ensuring compliance with various constraints and regulations. By focusing on responsible and ethical use, we aim to maximize the benefits of Q-Mamba while minimizing any adverse societal impacts.

### v2

@@
-Q-Mamba: On First Exploration of Vision Mamba for Image Quality Assessment
+QMamba: On First Exploration of Vision Mamba for Image Quality Assessment
 Abstract
-In this work, we take the first exploration of the recently popular foundation model, i.e., State Space Model/Mamba, in image quality assessment, aiming at observing and excavating the perception potential in vision Mamba.
+In this work, we take the first exploration of the recently popular foundation model, i.e., State Space Model/Mamba, in image quality assessment (IQA), aiming at observing and excavating the perception potential in vision Mamba.
 A series of works on Mamba has shown its significant potential in various fields, e.g., segmentation and classification.
-However, the perception capability of Mamba has been under-explored.
-Consequently, we propose Q-Mamba by revisiting and adapting the Mamba model for three crucial IQA tasks, i.e., task-specific, universal, and transferable IQA, which reveals that the Mamba model has obvious advantages compared with existing foundational models, e.g., Swin Transformer, ViT, and CNNs, in terms of perception and computational cost for IQA.
-To increase the transferability of Q-Mamba, we propose the StylePrompt tuning paradigm, where the basic lightweight mean and variance prompts are injected to assist the task-adaptive transfer learning of pre-trained Q-Mamba for different downstream IQA tasks.
-Compared with existing prompt tuning strategies, our proposed StylePrompt enables better perception transfer capability with less computational cost.
-Extensive experiments on multiple synthetic, authentic IQA datasets, and cross IQA datasets have demonstrated the effectiveness of our proposed Q-Mamba.
-footnotetext: ${}^{\textrm{\Letter}}$ Corresponding authors.
+However, the perception capability of Mamba remains under-explored.
+Consequently, we propose QMamba by revisiting and adapting the Mamba model for three crucial IQA tasks, i.e., task-specific, universal, and transferable IQA, which reveals its clear advantages over existing foundational models, e.g., Swin Transformer, ViT, and CNNs, in terms of perception and computational cost.
+To improve the transferability of QMamba, we propose the StylePrompt tuning paradigm, where lightweight mean and variance prompts are injected to assist task-adaptive transfer learning of pre-trained QMamba for different downstream IQA tasks.
+Compared with existing prompt tuning strategies, our StylePrompt enables better perceptual transfer with lower computational cost.
+Extensive experiments on multiple synthetic, authentic IQA datasets, and cross IQA datasets demonstrate the effectiveness of our proposed QMamba.
+The code will be available at: https://github.com/bingo-G/QMamba.git
+Machine Learning, ICML
+{guanfb, yuzihao, luyt31415}@mail.ustc.edu.cn, {xin.li, chenzhibo}@ustc.edu.cn
+Figure 1: Scanning Methodology Illustration.
+(a) VMamba’s scanning method(Liu et al., 2024c) flattens 2D data into 1D, impairing connectivity by distancing adjacent tokens.
+(b) Local scanning method scans within and across windows, placing semantically similar and distortion-related tokens closer, as shown in the blue boxes.
 1 Introduction
-Image Quality Assessment (IQA) aims to measure the subjective quality of images aligned with human perception, which has been applied in various visual fields, including visual acquisition, transmission, AIGC [25, 61, 63, 71], and UGC creation [58, 31, 42], etc.
-Establishing a great IQA metric is necessary to provide the right optimization direction tailored for image processing techniques, i.e., compression [30, 64, 8, 35], enhancement [27, 28, 24, 7], and ensure the perceptual quality of images.
-Early works on IQA have been achieved by leveraging the natural scene statistics in a hand-crafted manner [44, 46, 60, 52].
-With the advancements of deep neural networks (DNNs), learning-based IQA metrics [29, 38, 34, 76, 77, 3] have demonstrated significant potential for low-level perception, which can be roughly categorized into two types based on the pre-traiend backbones: CNN-based and Transformer-based methods.
-Although the impressive progress, learning-based IQA are susceptible to inherent limitations of existing pre-trained backbones: (i) the CNNs are skilled at learning local translation-invariant features from images while lacking enough long-range dependency modeling capability, hindering the global quality perception.
+Image Quality Assessment (IQA) aims to measure the subjective quality of images aligned with human perception, which has been applied in various visual fields, including visual acquisition, transmission, AIGC (Li et al., 2023a; Wang et al., 2023a), and UGC creation (Tu et al., 2021; Lu et al., 2024), etc.
+Establishing a great IQA metric is necessary to provide the right optimization direction tailored for image processing techniques, i.e., compression (Li et al., 2021; Wu et al., 2021; Yu et al., 2024b), enhancement (Li et al., 2023b; Fei et al., 2023), and ensure the perceptual quality of images.
+Early works on IQA have been achieved by leveraging the natural scene statistics in a hand-crafted manner (Mittal et al., 2012a, b).
+With the advancements of deep neural networks (DNNs), learning-based IQA metrics (Li et al., 2023c; Liu et al., 2022) have demonstrated significant potential for low-level perception, which can be roughly categorized into two types based on the pre-trained backbones: CNN-based and Transformer-based methods.
+Although the impressive progress, learning-based IQA is susceptible to inherent limitations of existing pre-trained backbones: (i) the CNNs are skilled at learning local translation-invariant features from images while lacking enough long-range dependency modeling capability, hindering the global quality perception.
 (ii) The emergence of Vision Transformers presents a great solution to model long-range dependency effectively by leveraging attention mechanisms.
 However, the quadratic complexity of self-attention operations poses unaffordable computational costs, especially for large-scale image quality assessment.
-Recently, an innovative foundation model, the State Space Model, particularly its implementation, i.e., Mamba [12] has shown considerable potential in various fields for balancing the computational costs and performances, e.g., segmentation[65, 43] and classification[54, 80].
-This raises one interesting question, i.e., “whether Mamba can surpass existing backbones on low-level visual perception", which is still under-explored.
-To answer this question, in this work, we initiate a new exploration of the Mamba model within the field of image quality assessment (IQA), and introduce the Q-Mamba, a newly developed IQA framework designed to address three key facets of IQA: task-specific, universal, and transferable image quality assessment.
+Recently, an innovative foundation model, the State Space Model, particularly its implementation, i.e., Mamba (Gu & Dao, 2023) has shown considerable potential in various fields for balancing the computational costs and performances, e.g., segmentation(Xing et al., 2024; Ma et al., 2024) and classification(Liu et al., 2024c; Zhu et al., 2024).
+This raises one interesting question, i.e., “whether Mamba can surpass existing backbones on low-level visual perception”, which is still under-explored.
+To answer this question, in this work, we initiate a new exploration of the Mamba model within the field of IQA and introduce the QMamba, a newly developed IQA framework designed to address three key facets of IQA: task-specific, universal, and transferable image quality assessment.
 Notably, since the lack of enough IQA dataset, learning-based IQA metrics entail the pre-trained backbone for perception knowledge extraction.
-Meanwhile, VMamba [39], as the most representative framework, has achieved excellent performance on high-level tasks by employing horizontal and vertical scanning strategies.
+Meanwhile, VMamba(Liu et al., 2024c), as the most representative framework, has achieved excellent performance on high-level tasks by employing horizontal and vertical scanning strategies.
 However, merely excavating the global perception knowledge is not optimal for IQA, since most artifacts affecting the image quality are related to local textures.
-Consequently, inspired by LocalMamba[18], we have adopted a local window scanning method, which significantly enhances Mamba’s ability to perceive local distortions, thereby demonstrating superior performance.
-We detail the architecture of Q-Mamba and examine different scanning methods to clarify how the model interacts with and processes image data.
-Our extensive analysis confirms that Q-Mamba surpasses traditional foundational models, showing superior perceptual accuracy and greater computational efficiency.
-Moreover, we have explored the perception transferability of Mamba across different datasets, i.e., different contents, and degradations.
-We can find that the Mamba-based IQA metric still suffers from severe performance drop when they encounter large domain shifts between synthetic distortions[53, 33], authentic distortions[10, 17], and Artificial Intelligence-Generated Content (AIGC)[25, 61] distortions, etc.
-To further amplify the transferability of Q-Mamba across a range of IQA applications, we introduce a simple but effective tuning strategy called StylePrompt.
-This is based on the finding that the domain shifts in IQA tend to correlate with their feature statistics/style [41], such as mean and variance.
-Concretely, our StylePrompt aims to adaptively adjust the mean and variance of pre-trained Q-Mamba towards the target IQA tasks by setting a group of light-weight learnable $1\times 1\times C$ parameters.
-Extensive experiments have shown that our StylePrompt greatly facilitates the task-adaptive learning of Q-Mamba, enabling efficient knowledge transfer across diverse IQA tasks with fewer parameter costs.
+Consequently, inspired by LocalMamba(Huang et al., 2024), we have adopted a local window scanning method, which significantly enhances ability of Mamba to perceive local distortions, thereby demonstrating superior performance.
+We detail the architecture of QMamba and examine different scanning methods to clarify how the model interacts with and processes image data.
+Our extensive analysis confirms that QMamba surpasses traditional foundational models, showing superior perceptual accuracy and greater computational efficiency.
+Moreover, we have explored the perceptual transferability of Mamba across different datasets, i.e., different contents, and degradations.
+We can find that the Mamba-based IQA metric still suffers from severe performance drop when they encounter large domain shifts between synthetic distortions(Sheikh et al., 2006; Lin et al., 2019), authentic distortions(Ghadiyaram & Bovik, 2015; Hosu et al., 2020), and Artificial Intelligence-Generated Content (AIGC)(Li et al., 2023a; Wang et al., 2023a) distortions, etc.
+To further amplify the transferability of QMamba across a range of IQA applications, we introduce a simple but effective tuning strategy called StylePrompt.
+This is based on the finding that the domain shifts in IQA tend to correlate with their feature statistics/style (Lu et al., 2022), such as mean and variance.
+Concretely, our StylePrompt aims to adaptively adjust the mean and variance of pre-trained QMamba towards the target IQA tasks by setting a group of light-weight learnable $1\times 1\times C$ parameters.
+Extensive experiments have shown that our StylePrompt greatly facilitates the task-adaptive learning of QMamba, enabling efficient knowledge transfer across diverse IQA tasks with fewer parameter costs.
 The main contributions of this paper are summarized as follows:
-  • We embark on a novel exploration of the Mamba model within image quality assessment (IQA), and propose the Q-Mamba, a powerful IQA metric for three critical tasks of IQA: task-specific, universal, and transferable image quality assessment.
+  • We embark on a novel exploration of the Mamba model within image quality assessment, and propose the QMamba, a powerful IQA metric for three critical tasks of IQA: task-specific, universal, and transferable image quality assessment.
 This exploration has demonstrated the superior potential of Mamba for subjective perception, advancing the development of IQA.
-  • To improve the perception transferability of the Q-Mamba, we introduce a simple but effective tuning strategy, i.e., StylePrompt.
-This strategy enable the efficient knowledge transfer of pre-trained Q-Mamba for downstream IQA tasks, while only tuning fewer learnable parameters to adjust the statistics of perception features.
-  • Extensive experiments have shown that our Q-Mamba has consistently achieved state-of-the-art results on various prominent IQA datasets compared with existing IQA methods, thereby validating the efficacy of the Mamba model in quality assessment.
+  • To improve the perception transferability of the QMamba, we introduce a simple but effective tuning strategy, i.e., StylePrompt.
+This strategy enables the efficient knowledge transfer of pre-trained QMamba for downstream IQA tasks, while only tuning fewer learnable parameters to adjust the statistics of perception features.
+  • Extensive experiments have shown that our QMamba has consistently achieved state-of-the-art results on various prominent IQA datasets compared with existing IQA methods, thereby validating the efficacy of the Mamba model in quality assessment.
 Moreover, our StylePrompt achieves nearly equivalent performance to full model tuning while utilizing only $4\%$ of the whole parameters, demonstrating its effectiveness in perception knowledge transfer scenarios.
@@
 2.1 Blind Image Quality Assessment (BIQA)
-Early BIQA methods relied heavily on manually designed features for quality score regression [44, 46, 60, 52].
+Early BIQA methods relied heavily on manually designed features for quality score regression (Mittal et al., 2012a, b; Venkatanath et al., 2015; Saad et al., 2012; Min et al., 2018).
 However, these handcrafted features were insufficient for addressing the complexity of BIQA tasks.
@@
 CNNs have demonstrated robust feature extraction capabilities, leading to their widespread adoption in BIQA tasks.
-Early works like CNNIQA [20] used convolutional models for feature learning and quality regression, substantially outperforming handcrafted features.
-DBCNN [75] introduced a dual-stream network to address synthetic and authentic distortions separately, integrating these insights for better quality prediction.
-NIMA [57] and PQR [72] leveraged pre-trained models on ImageNet for quality score prediction, enhancing accuracy through well-established neural architectures.
-MetaIQA [79] employed meta-learning to adapt to unknown distortions by learning shared priors for various distortion types, while HyperIQA [56] used a hypernetwork to adaptively establish perceptual rules, improving generalization.
-Despite these advancements, CNNs’ local bias limits their ability to fully exploit both global and local information in BIQA tasks.
+Early works like CNNIQA (Kang et al., 2014) used convolutional models for feature learning and quality regression, substantially outperforming handcrafted features.
+DBCNN (Zhang et al., 2020) introduced a dual-stream network to address synthetic and authentic distortions separately, integrating these insights for better quality prediction.
+NIMA (Talebi & Milanfar, 2018) and PQR (Zeng et al., 2017) leveraged pre-trained models on ImageNet for quality score prediction, enhancing accuracy through well-established neural architectures.
+MetaIQA (Zhu et al., 2020) employed meta-learning to adapt to unknown distortions by learning shared priors for various distortion types, while HyperIQA (Su et al., 2020) used a hypernetwork to adaptively establish perceptual rules, improving generalization.
+Despite these advancements(Saha et al., 2023; Zhao et al., 2023; Chen et al., 2024a), CNNs’ local bias limits their ability to fully exploit both global and local information in BIQA tasks.
+Figure 2: Architectural Overview of QMamba Framework and the Detailed of StylePrompt Tuning Mechanism.
 Transformer-based BIQA.
 Transformers offer superior global modeling capabilities compared to CNNs.
-TReS [11] addressed CNNs’ local bias by capturing local structural information with CNNs and then using Transformers for sequential feature extraction.
-MUSIQ [21] designed a multi-scale image Transformer architecture capable of handling images with varying sizes and aspect ratios.
-DEIQT [50] leveraged a Transformer-based BIQA architecture with attention mechanisms to align with human perception, enhancing model performance and reducing prediction uncertainty.
+TReS (Golestaneh et al., 2022) addressed CNNs’ local bias by capturing local structural information with CNNs and then using Transformers for sequential feature extraction.
+MUSIQ (Ke et al., 2021) designed a multi-scale image Transformer architecture capable of handling images with varying sizes and aspect ratios.
+DEIQT (Qin et al., 2023) leveraged a Transformer-based BIQA architecture with attention mechanisms to align with human perception, enhancing model performance and reducing prediction uncertainty.
 However, the quadratic complexity of Transformers presents a challenge, highlighting the need for architectures with linear complexity capable of global modeling for BIQA tasks.
+In addition to the above methods, many recent works (Zhang et al., 2023; Shin et al., 2024; Yu et al., 2024a) have also adopted Transformer-based designs and achieved competitive performance, yet they still face limitations due to high computational complexity.
 2.2 State Space Models (SSMs)
-State-space models, known for their linear complexity in capturing long-range dependencies, have been integrated into deep learning architectures.
-The Structured State-Space model (S4) [14] was a pioneer in deep state-space modeling for remote dependency modeling.
-Subsequent advancements [13, 55, 9] further propelled the development of state-space models.
-Mamba [12], by integrating selection mechanisms and hardware-aware algorithms, has shown effective long-range modeling capabilities with linear complexity growth.
+State space models, known for their linear complexity in capturing long-range dependencies, have been integrated into deep learning architectures.
+The Structured State Space model (S4) (Gu et al., 2021) was a pioneer in deep state space modeling for remote dependency modeling.
+Subsequent advancements (Gu et al., 2020; Smith et al., 2022; Fu et al., 2022) further propelled the development of state space models.
+Mamba (Gu & Dao, 2023), by integrating selection mechanisms and hardware-aware algorithms, has shown effective long-range modeling capabilities with linear complexity growth.
 Initially focused on NLP tasks, Mamba has rapidly expanded into other domains.
-Vim [80] introduced a bidirectional SSM block for visual representation learning, achieving performance comparable to ViT [5].
-VMamba [39] introduced a cross-scan module to traverse spatial domains and convert non-causal visual images into ordered block sequences, maintaining linear complexity while retaining global receptive fields.
-LocalMamba [18] employed a window-based scanning approach to integrate local inductive biases, enhancing the visual Mamba model.
-These advancements validate Mamba’s efficacy in visual tasks, leading to its application in image classification [39, 80, 47, 67, 1], video understanding [62, 4, 26], image restoration [15, 54, 78], point cloud analysis [32, 74, 37], and biomedical image segmentation [43, 65, 51, 36].
-This paper explores Mamba’s potential in visual quality perception tasks, providing a simple yet effective baseline for future quality perception research.
+Vim (Zhu et al., 2024) introduced a bidirectional SSM block for visual representation learning, achieving performance comparable to ViT (Dosovitskiy et al., 2020).
+VMamba (Liu et al., 2024c) introduced a cross-scan module to traverse spatial domains and convert non-causal visual images into ordered block sequences, maintaining linear complexity while retaining global receptive fields.
+LocalMamba (Huang et al., 2024) employed a window-based scanning approach to integrate local inductive biases, enhancing the visual Mamba model.
+These advancements validate the efficacy of Mamba in visual tasks, leading to its application in image classification (Liu et al., 2024c; Zhu et al., 2024; Patro & Agneeswaran, 2024), video understanding (Wang et al., 2023b; Chen et al., 2024b; Li et al., 2024), image restoration (Guo et al., 2024; Shi et al., 2024; Zhen et al., 2024), point cloud analysis (Liang et al., 2024; Zhang et al., 2024; Liu et al., 2024b), and biomedical image segmentation (Ma et al., 2024; Xing et al., 2024; Liu et al., 2024a).
+These studies have demonstrated the effectiveness of state space models in visual tasks, providing a solid foundation for further exploration of their potential in visual perception.
 3 Method
-3.1 Prelimineries
-State Space Model (SSM).
-Traditional State Space Models (SSMs) utilize first-order differential equations to establish a linear time-invariant system, translating one-dimensional input sequences $x(t)\in\mathbb{R}$ into outputs through hidden states $h(t)$ and $y(t)$, described by:
-$\begin{gathered}h^{\prime}(t)=Ah(t)+Bx(t)\\ y(t)=Ch(t)\end{gathered}$ (1)
-For compatibility with deep learning’s sequential input requirements, these models undergo a transition to a discrete framework within the S4 model, using discretization parameters $\Delta$ to modify $A$ and $B$ into $\overline{A}$ and $\overline{B}$:
-$\begin{gathered}h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t}\\ y_{t}=Ch_{t}\end{gathered}$ (2)
-$\begin{gathered}\overline{A}=\exp(\Delta A)\\ \overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B\end{gathered}$ (3)
-Mamba[12], evolving from these foundations, introduces a selection algorithm for dynamic adaptation of $A$, $B$, and $C$ based on input, enhancing model response to complex sequences and contextual variations.
-However, this necessitates abandoning traditional convolutional methods for kernel fusion and parallel scanning to maintain computational efficiency, effectively scaling with sequence length and providing advanced capabilities in sequence generation for language models.
-Figure 1: Overview of the Q-Mamba architecture and description of the StylePrompt.
 3.1 Exploring Mamba for Perception
 3.1.1 Overall Framework
-The overall architecture of our Q-Mamba model is illustrated in Figure1.
-To successfully incorporate the Mamba model into visual perception tasks, we implemented the Q-Mamba architecture following the construction method of the VMamba’s visual Mamba blocks.
-This architecture consists of multiple layers of VMamba blocks, with the SS2D block serving as the core module, which facilitates information transmission through a residual network.
-The residual network incorporates convolutional layers for feature extraction and activation layers for computing gating signals.
-We employed multiple network stages, each comprising a downsampling layer and a VMamba Block, which construct multi-level representations at varying resolutions through downsampling, thus extracting richer perceptual features.
-To explore the impact of model size on quality perception tasks, we experimented with three different scales of the model: QMamba-Tiny, QMamba-Small, and QMamba-Base.
+The overall architecture of our proposed QMamba model is depicted in Figure 2.
+To address the unique challenges of visual perception tasks, we developed a novel architecture that rethinks the design principles of state space models for visual data.
+The network incorporates a hierarchical residual structure, where convolutional layers enable effective feature extraction while specialized activation layers compute adaptive gating signals.
+Our architecture organizes the processing into multiple network stages, each combining a strategic downsampling layer with our enhanced Mamba-based processing block.
+This design enables the construction of multi-level representations at varying resolutions, facilitating the extraction of richer perceptual features through progressive abstraction.
+To systematically investigate the relationship between model capacity and quality perception performance, we developed three distinct variants of our architecture: QMamba-Tiny, QMamba-Small, and QMamba-Base.
+Each variant maintains the core architectural principles while scaling in complexity, allowing us to explore the trade-offs between computational efficiency and perceptual accuracy across different application scenarios.
 3.1.2 Perception with Local Scanning
-Originally, Mamba demonstrated exceptional proficiency in NLP tasks, characterized by inherently causal sequential inputs.
-However, when adapted to visual tasks, which lack corresponding causal frameworks, the model encountered substantial difficulties.
-It struggled to effectively discern the complex interdependencies among image pixels, thereby impeding its learning processes.
-VMamba[39] employs a novel bidirectional horizontal and vertical scanning mechanism, illustrated in Figure 2(a), to convert two-dimensional data into one-dimensional inputs for model ingestion.
-Although this method facilitates global modeling of pixel sequences, it unfortunately causes locally adjacent tokens to be spaced significantly apart, thus failing to capture essential local dependencies.
-This limitation becomes particularly acute in quality assessment tasks, where accurate distortion detection is dependent on the granularity of local detail.
-Drawing inspiration from LocalMamba[18], we have implemented a window-based scanning strategy, as depicted in Figure 2(b), to enhance local perceptual accuracy.
-This involves conducting horizontal scans within each window followed by comprehensive window-level horizontal scans, with a parallel approach adopted in the vertical dimension.
-This method markedly enhances VMamba’s capability to detect local quality details.
-Despite the model’s input sequence undergoing downsampling and aggregation, this scanning approach retains an ability for global perception, ensuring a robust analytical balance that effectively captures both detailed local nuances and broader contextual information.
-Figure 2: Scanning Methodology Illustration.
-(a) VMamba’s[39] method flattens 2D data into 1D, impairing connectivity by distancing adjacent tokens.
-(b) Our local scanning method[18] scans within and across windows, placing semantically similar and distortion-related tokens closer, as shown in the blue boxes.
+While the original Mamba model excels in natural language processing tasks with inherently causal inputs, it faces challenges when applied to visual tasks due to the absence of spatial causality.
+In particular, it struggles to capture complex spatial dependencies among image pixels, which hinders its ability to model local distortions effectively.
+To address this, VMamba (Liu et al., 2024c) introduces a bidirectional horizontal and vertical scanning strategy to convert 2D images into 1D sequences suitable for sequential modeling.
+Although this enables global pixel-level modeling, it disrupts the continuity of locally adjacent tokens, weakening the model’s ability to perceive fine-grained distortions—an essential factor for image quality assessment.
+Inspired by LocalMamba (Huang et al., 2024), we adopt a window-based scanning approach that performs horizontal scans within local windows, followed by window-level scans, and applies the same strategy vertically.
+This method enhances local distortion perception while retaining global awareness through hierarchical composition.
+Despite downsampling and aggregation in the input, this scanning scheme maintains a balance between local detail and broader context.
+Unlike LocalMamba, which relies on attention-based dynamic routing and suffers from unstable inference and high computational cost, LQMamba adopts a hierarchical architecture with fixed-size windows that change with network depth.
+As illustrated in Figure 1 (b), this structure enables the model to capture multi-scale perceptual cues, from fine-grained distortions to broader contextual patterns, while maintaining stable and efficient inference.
+It achieves a good balance between accuracy and efficiency, making it well-suited for IQA tasks that demand consistent perception across various distortion types and spatial scales.
+Figure 3: t-SNE Visualization of Distortion-Specific Feature Separation: QMamba vs. Conventional Backbones
+3.1.3 Analysis of Perceptual Capability in State Space Models
+To investigate the feature selection mechanism of State Space Models (i.e.,Mamba) in visual quality assessment tasks, we employ t-SNE visualization for deep-layer feature analysis, as illustrated in Figure 3.
+Visualization results reveal that Mamba exhibits distinctive characteristic evolution patterns through dynamic state updating mechanisms, progressively enhancing distortion semantic perception across network hierarchies.
+Specifically, shallow layers preserve diverse features of original visual signals, while deeper layers adaptively filter redundant background information through gated state selection, intensifying distortion-type-specific feature focusing.
+This drives distortion-homogeneous samples to form cluster structures in feature space while amplifying inter-class discriminability.
+Such feature refinement strategy not only retains critical discriminative information for quality assessment but also significantly improves model sensitivity to quality degradation cues, thereby strengthening prediction robustness.
 3.2 Tuning the Mamba with StylePrompt
-Although the Mamba architecture offers reduced computational complexity compared to other models, achieving higher performance still necessitates a substantial number of parameters, which poses a challenge for efficient and transferable learning critical in Image Quality Assessment (IQA) tasks.
-Prompt tuning[19, 23], a common method for efficient fine-tuning, presents a promising approach.
-However, conventional prompt tuning techniques involving convolutional or attention operations still require significant parameter overhead[49].
-Considering the need for transfer learning in IQA tasks to adapt specifically to the current distortion style, we propose the StylePrompt to enhance the transferability of IQA capabilities.
-Our method achieves results comparable to full-parameter fine-tuning with a minimal parameter increase.We illustrate the StylePrompt we used in Figure1.
-Our StylePrompt consists of two parts, which are described in detail below:
+Although the Mamba architecture reduces computational complexity compared to other models, achieving higher performance still requires a substantial number of parameters, which poses a challenge for efficient transferable learning, crucial for IQA tasks.
+We observe that domain shifts in IQA tasks tend to correlate with their feature statistics or style (Lu et al., 2022), such as mean and variance.
+Building on this insight, we propose a lightweight tuning strategy, StylePrompt, designed to adjust the mean and variance of the pre-trained QMamba features, thereby aligning them with the distortions and content types of the target domain.
+This approach enables us to achieve results comparable to full-parameter fine-tuning while using a minimal number of parameters, significantly enhancing both the efficiency and performance of QMamba in transferable IQA tasks.
+Figure 2 illustrates the StylePrompt, which consists of two components that will be described in detail below:
 3.2.1 StylePrompt Generation (SPG)
+We designed the StylePrompt Generation phase to facilitate the creation of prompts and their interaction with the original features.
 In a multi-stage network architecture, as images progress through each stage of the network, we learn a set of prompts $P_{s}\in\mathbb{R}^{N\!\times\!1\!\times\!1\!\times\!C}$, containing N prompt components designed to generate affine parameters for style adaptation.
-These prompts are specifically utilized to inject the distortion style of the current image into the features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, facilitating learning of the style pertinent to the current domain.
-We have devised a straightforward StylePrompt Generation stage where the prompt components $P_{s}$ are initially formed as a set of learnable parameters, interacting with the incoming features to embed distortion information.
+These prompts are specifically utilized to inject the distortion information of the current data into the features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, facilitating learning of the style pertinent to the target domain.
 To enable the prompt components to extract specific style information from the input stream dynamically, we predict the weights for different prompt components based on the input features.
@@
 These weights are then applied to the multiple prompt components to amalgamate them into a new prompt $P_{f}$, effectively encapsulating the current style information.
-The operation can be succinctly summarized by the following formula:
+The operation can be briefly summarized by the following formula:
 $P_{f}=\sum_{c=1}^{N}w_{s}P_{s},\quad w_{s}=\text{Softmax}(\text{Conv1x1}(\text{GAP}(F_{i})))$ (1)
 3.2.2 StylePrompt Injection (SPI)
-In the StylePrompt Generation phase, we created the fused prompt $P_{f}$, which contains the distortion style information of the current domain.
-During the StylePrompt Injection process, this style information is injected into the original features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$ produced by the current layer.
-We adapt the distribution of the original features to the style by learning affine parameters.
-Specifically, as each layer utilizes a fixed-size prompt component, the fused Prompt generated in the StylePrompt Generation phase is matched to the channel dimensions of the current features through linear layers designed to align dimensions.
-This vector is then processed through two specific linear layers whose primary function is to adjust dimensions, thereby producing the affine parameters $\gamma_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$ and $\beta_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$.
-These parameters are used to inject the distorted style information solely on the channel dimension into the original feature distribution.
+In the SPG phase, the fused prompt $P_{f}$ is created, containing the distortion style information of the target domain.
+During the subsequent StylePrompt Injection process, this style information is injected into the original features $F_{i}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$ produced by the current layer.
+To achieve this, $P_{f}$ is matched to the channel dimensions of the current features using linear layers designed for dimensional alignment.
+The processed prompt then generates affine parameters $\gamma_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$ and $\beta_{v}\in\mathbb{R}^{1\times 1\times\hat{C}}$, which modulate the mean and variance of the original feature distribution solely along the channel dimension.
+This adaptive adjustment ensures that the feature distribution is tailored to the distortion style of the target domain with minimal computational overhead.
 The process of our StylePrompt Injection can be summarized as follows:
@@
 $\beta_{v}=\text{Linear}_{\beta}(\text{Conv}(P_{f}))$ (3)
-$F{{}^{\prime}}_{i}=F_{i}\cdot(1+\gamma)+\beta$ (7)
+$F{{}^{\prime}}_{i}=F_{i}\cdot(1+\gamma_{v})+\beta_{v}$ (4)
 Specifically, $F_{i}^{\prime}$ represents the original features after the injection of style information. These enhanced features will serve as the new input to the subsequent layer of the network.
-4 Experiments
-4.1 Experimental Setup
-4.1.1 Datasets
-We conducted foundational experiments on ten popular IQA datasets, which include four synthetic datasets: LIVE[53], CSIQ[22], TID2013[48], and KADID[33]; four authentic datasets: LIVEC[10], KonIQ[17], LIVEFB[68], and SPAQ[6]; and two AIGC datasets: AIGC2023[61] and AGIQA3K[25].
-In our single-task experiments, we utilized all ten datasets for testing.
-We report the results for the eight datasets consisting of synthetic and authentic datasets in the main text, including comparisons with existing SOTA BIQA methods and performance comparisons of Swin Transformer, ViT, and ResNet across different parameter versions.
-In the appendix, we report the comparative results of the two AIGC datasets against the aforementioned model architectures.
-In our universal IQA and transferable IQA experiments, we selected six datasets for training and testing, which include two synthetic datasets (LIVE[53], KADID[33]), two authentic datasets (LIVEC[10], KonIQ[17]), and two AIGC datasets (AIGC2023[61], AGIQA3K[25]).
-4.1.2 Evaluation Criteria
-The evaluation metrics employed in our study are the widely utilized Pearson Linear Correlation Coefficient (PLCC) and Spearman’s Rank Correlation Coefficient (SRCC), both of which range from 0 to 1.
-Values approaching 1 denote a higher degree of prediction relevance.
-4.1.3 Experimental Details
-Our experimental methodology closely adheres to the training strategy outlined in DEIQT[50], where input images are randomly cropped into ten patches of 224×224 resolution.
-We utilized three variants of the VMamba architecture: VMamba-B, VMamba-S, and VMamba-T.
-Both VMamba-B and VMamba-S feature an encoder with a depth of 15 layers, where VMamba-B incorporates an embedding dimension of 128, and VMamba-S has an embedding dimension of 96.
-In contrast, VMamba-T is characterized by a reduced depth of 4 layers and an embedding dimension of 96.
-Training procedures leveraged weights pre-trained on the ImageNet-1K dataset, spanning across a total of 9 epochs, with batch sizes adjusted according to the respective dataset sizes.
-All experiments were conducted utilizing multiple NVIDIA RTX 4090 GPUs.
-Table 1: Performance comparison for task-specific IQA. Bold indicates the top two results.
-  Model LIVE CSIQ TID2013 KADID LIVEC KonIQ LIVEFB SPAQ
-  FLOPs PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
-     ILNIQE[73] - 0.906 0.902 0.865 0.822 0.648 0.521 0.558 0.534 0.508 0.508 0.537 0.523 0.332 0.294 0.712 0.713 0.618
-BRISQUE[45] - 0.944 0.929 0.748 0.812 0.571 0.626 0.567 0.528 0.629 0.629 0.685 0.681 0.341 0.303 0.817 0.809 0.664
-WaDIQaM[2] - 0.955 0.960 0.844 0.852 0.855 0.835 0.752 0.739 0.671 0.682 0.807 0.804 0.467 0.455 - - 0.763
-DBCNN[75] - 0.971 0.968 0.959 0.946 0.865 0.816 0.856 0.851 0.869 0.851 0.884 0.875 0.551 0.545 0.915 0.911 0.852
-TIQA[70] - 0.965 0.949 0.838 0.825 0.858 0.846 0.855 0.850 0.861 0.845 0.903 0.892 0.581 0.541 - - 0.829
-MetaIQA[79] - 0.959 0.960 0.908 0.899 0.868 0.856 0.775 0.762 0.802 0.835 0.856 0.887 0.507 0.540 - - 0.815
-P2P-BM[69] - 0.958 0.959 0.902 0.899 0.856 0.862 0.849 0.840 0.842 0.844 0.885 0.872 0.598 0.526 - - 0.835
-HyperIQA[56] - 0.966 0.962 0.942 0.923 0.858 0.840 0.845 0.852 0.882 0.859 0.917 0.906 0.602 0.544 0.915 0.911 0.858
-TReS[11] - 0.968 0.969 0.942 0.922 0.883 0.863 0.858 0.859 0.877 0.846 0.928 0.915 0.625 0.554 - - 0.858
-MUSIQ[21] - 0.911 0.940 0.893 0.871 0.815 0.773 0.872 0.875 0.746 0.702 0.928 0.916 0.661 0.566 0.921 0.918 0.832
-DEIQT[50] - 0.982 0.980 0.963 0.946 0.908 0.892 0.887 0.889 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919 0.884
-LoDa[66] - 0.979 0.975 - - 0.901 0.869 0.936 0.931 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925 0.882
+  LIVE CSIQ TID2013 KADID LIVEC KonIQ LIVEFB SPAQ
+    Method GFLOPS PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
+    ILNIQE - 0.906 0.902 0.865 0.822 0.648 0.521 0.558 0.534 0.508 0.508 0.537 0.523 0.332 0.294 0.712 0.713 0.618
+BRISQUE - 0.944 0.929 0.748 0.812 0.571 0.626 0.567 0.528 0.629 0.629 0.685 0.681 0.341 0.303 0.817 0.809 0.664
+WaDIQaM(5.24M) 2.43G 0.955 0.960 0.844 0.852 0.855 0.835 0.752 0.739 0.671 0.682 0.807 0.804 0.467 0.455 - - 0.763
+DBCNN(15.31M) 16.51G 0.971 0.968 0.959 0.946 0.865 0.816 0.856 0.851 0.869 0.851 0.884 0.875 0.551 0.545 0.915 0.911 0.852
+TIQA(23.68M) - 0.965 0.949 0.838 0.825 0.858 0.846 0.855 0.850 0.861 0.845 0.903 0.892 0.581 0.541 - - 0.829
+MetaIQA(13.24M) 1.82G 0.959 0.960 0.908 0.899 0.868 0.856 0.775 0.762 0.802 0.835 0.856 0.887 0.507 0.540 - - 0.815
+HyperIQA(27.38M) 4.31G 0.966 0.962 0.942 0.923 0.858 0.840 0.845 0.852 0.882 0.859 0.917 0.906 0.602 0.544 0.915 0.911 0.858
+TReS(152.45M) 20.03G 0.968 0.969 0.942 0.922 0.883 0.863 0.858 0.859 0.877 0.846 0.928 0.915 0.625 0.554 - - 0.858
+MUSIQ(27.13M) 9.02G 0.911 0.940 0.893 0.871 0.815 0.773 0.872 0.875 0.746 0.702 0.928 0.916 0.661 0.566 0.921 0.918 0.832
+DEIQT(24.04M) 5.41G 0.982 0.980 0.963 0.946 0.908 0.892 0.887 0.889 0.894 0.875 0.934 0.921 0.663 0.571 0.923 0.919 0.884
+LoDa(*) 23.74G 0.979 0.975 - - 0.901 0.869 0.936 0.931 0.899 0.876 0.944 0.932 0.679 0.578 0.928 0.925 0.882
     ResNet-50(23.51M) 4.11G 0.879 0.884 0.861 0.841 0.747 0.686 0.784 0.786 0.868 0.831 0.908 0.886 0.313 0.269 0.907 0.907 0.772
@@
 LQMamba-B(93.79M) 15.30G 0.959 0.951 0.915 0.889 0.965 0.964 0.943 0.941 0.913 0.888 0.947 0.933 0.675 0.582 0.934 0.929 0.896
+    LoDa has a total of 118.23M model parameters and 8.93M trainable parameters.
+Table 1: Performance Comparison for Task-Specific IQA. Bold Indicates the Top Two Results.
+4 Experiments
+4.1 Experimental Setup
+4.1.1 Datasets
+We conducted foundational experiments on ten popular IQA datasets, which include four synthetic datasets: LIVE(Sheikh et al., 2006), CSIQ(Larson & Chandler, 2010), TID2013(Ponomarenko et al., 2015), and KADID(Lin et al., 2019); four authentic datasets: LIVEC(Ghadiyaram & Bovik, 2015), KonIQ(Hosu et al., 2020), LIVEFB(Ying et al., 2020), and SPAQ(Fang et al., 2020); and two AIGC datasets: AIGC2023(Wang et al., 2023a) and AGIQA3K(Li et al., 2023a).
+We will provide a more detailed introduction to these datasets in Appendix.
+4.1.2 Evaluation Criteria
+The evaluation metrics employed in our study are the widely utilized Pearson Linear Correlation Coefficient (PLCC) and Spearman’s Rank Correlation Coefficient (SRCC), both of which range from 0 to 1.
+Values approaching 1 denote a higher degree of prediction relevance.
+4.1.3 Experimental Details
+Our experimental methodology closely follows the training strategy outlined in DEIQT(Qin et al., 2023), where input images are randomly cropped into ten patches, each with a resolution of 224×224.
+We employed three variants of the VMamba architecture: QMamba-B, QMamba-S, and QMamba-T.
+Both QMamba-B and QMamba-S feature an encoder with a depth of 15 blocks, with QMamba-B incorporating an embedding dimension of 128, and QMamba-S using an embedding dimension of 96.
+In contrast, QMamba-T is designed with a reduced depth of 4 blocks and an embedding dimension of 96.
+Training procedures leveraged weights pre-trained on the ImageNet-1K dataset, spanning a total of 9 epochs.
+Batch sizes were adjusted according to the respective dataset sizes, e.g., 32 for LIVEC and 128 for KonIQ.
+We used the AdamW optimizer for training, with the learning rate set to $2\times 10^{-4}$ and a decay factor of 10 applied every 3 epochs.
+We compared the performance of ResNet(He et al., 2016), ViT(Dosovitskiy et al., 2020), and Swin Transformer(Liu et al., 2021), all of which were implemented using the official versions and loaded with pre-trained weights.
+To ensure fairness, other experimental settings were kept as consistent as possible.
+All experiments were conducted using multiple NVIDIA RTX 4090 GPUs.
 4.2 A Comparison Between Different IQA Backbones
@@
 We conducted comprehensive training and testing across the ten datasets previously introduced, drawing analytical comparisons based on results reported by existing methods.
-Given the predominance of existing BIQA methods targeting synthetic and distorted datasets, we present in Table 1 a detailed comparative analysis of the performance of state-of-the-art (SOTA) methods alongside popular architectures such as ResNet[16], ViT[5], and Swin Transformer[40], against our Q-Mamba architecture.
+For this task, $80\%$ of the images in each dataset were used for training, while the remaining $20\%$ were reserved for testing.
+Given the predominance of BIQA methods targeting synthetic and authentic datasets, Table 1 presents a detailed comparative analysis of state-of-the-art (SOTA) methods and popular architectures such as ResNet, ViT, and Swin Transformer, in comparison to our QMamba architecture.
 This comparison elucidates performance discrepancies and computational complexities across different parameter configurations.
-Our findings demonstrate that the LocalQMamba configuration achieves optimal performance across six datasets and, when juxtaposed with other models of comparable parameter scales, it exhibits lower Gflops, thereby confirming the efficacy and reduced computational complexity of the Mamba architecture.
-Comparative results for the two AIGC datasets are documented in the appendix in Table 7.
+The results in the table demonstrate that the LQMamba configuration achieves optimal performance across six datasets.
+Compared to existing IQA methods with similar parameter counts, QMamba-T exhibits lower GFLOPS and superior performance, confirming the efficiency and reduced computational complexity of the Mamba architecture.
+Comparative results for the two AIGC datasets are documented in the appendix.
+                 Mixed Training
+                 LIVE & KADID & LIVEC & KonIQ & AGIQA3K & AIGCIQA2023
+     Method Parameters GFLOPS PLCC_Average SRCC_Average
+     ResNet-50 23.51M 4.11G 0.878 0.853
+ViT-S 21.67M 4.61G 0.891 0.867
+Swin-T 27.52M 4.51G 0.900 0.883
+     DEIQT 24.04M 5.41G 0.895 0.873
+LoDa 8.93M* 23.68G 0.876 0.855
+     QMamba-T 27.99M 4.47G 0.905 0.886
+LQMamba-T 29.87M 4.44G 0.909 0.888
+  • *Trainable parameters.
+Table 2: Performance Comparison for Universal IQA.
+  Train KonIQ & LIVEC (Authentic) KADID & LIVE (Synthetic)
+  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K
+    Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
+    DEIQT 0.583 0.595 0.558 0.564 0.802 0.794 0.698 0.632 0.518 0.526 0.579 0.532 0.580 0.575 0.512 0.506 0.601
+LoDa 0.600 0.600 0.733 0.729 0.808 0.800 0.715 0.650 0.527 0.513 0.558 0.499 0.592 0.586 0.564 0.541 0.622
+Without_tuning 0.535 0.501 0.726 0.744 0.789 0.784 0.726 0.744 0.553 0.549 0.595 0.550 0.600 0.612 0.675 0.649 0.642
+Lin_Probe(R) 0.606 0.581 0.783 0.811 0.824 0.804 0.769 0.695 0.871 0.848 0.785 0.751 0.803 0.788 0.771 0.723 0.755
+Full_tuning (93.79M)
+0.936 0.930 0.941 0.940 0.885 0.863 0.910 0.851 0.943 0.928 0.899 0.873 0.880 0.859 0.910 0.856 0.908
+StylePrompt (3.83M)
+0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.901
+StylePrompt & R 0.921 0.913 0.944 0.945 0.877 0.853 0.906 0.847 0.931 0.912 0.889 0.860 0.876 0.857 0.905 0.849 0.898
+Table 3: Performance Comparison for Transferable IQA.
 4.2.2 Universal IQA
-The efficacy of the Q-Mamba architecture was rigorously assessed across a spectrum of datasets, adopting a joint training protocol on six varied datasets, encompassing two synthetic datasets (LIVE, KADID), two authentic datasets (LIVEC, KonIQ), and two AIGC datasets (AIGC2023, AGIQA3K).
-Subsequent to this joint training phase, we evaluated the performance outcomes, with the findings depicted in Table 2.
-These results highlight Q-Mamba’s robust multi-tasking capabilities.
-In comparison to several established foundational models, Q-Mamba consistently delivered notable performance across the majority of the datasets, thereby substantiating its versatility and effectiveness in managing generalized tasks with high efficacy.
-Figure 3: t-SNE clustering of different backbones for various types of distortions
+We evaluated the effectiveness of the QMamba architecture for universal tasks by employing mixed training across six different datasets: two synthetic datasets (LIVE, KADID), two authentic datasets (LIVEC, KonIQ), and two AIGC datasets (AIGC2023, AGIQA3k).
+For each dataset, $20\%$ of the data was reserved for performance testing.
+The average results of models with similar scales are presented in Table 2, with detailed results provided in Appendix.
+These findings highlight strong multi-tasking capabilities of Mamba.
+Compared to several mainstream models, QMamba performed well across most datasets, confirming its effectiveness in handling general tasks.
 4.2.3 Analysis
-To elucidate the reasons behind the VMamba architecture’s notable performance on task-specific IQA and universal IQA tasks, we engaged a subset of the KADID dataset featuring seven distinct types of distortions.
-We conducted t-SNE[59] dimensionality reduction clustering on the features processed through four disparate model architectures, as illustrated in Figure 3.
-The clustering outcomes demonstrate that the VMamba architecture exhibits exceptional distortion perception capabilities, effectively distinguishing various types of distortions and surpassing the performance of the ViT structure.
-Conversely, the ResNet and Swin Transformer architectures did not manifest significant abilities in aggregating different distortion types.
-Consequently, it becomes apparent that VMamba’s exemplary performance across diverse quality assessment tasks can be attributed, in part, to its robust distortion perception capabilities.
-Table 2: Performance comparison for universal IQA. Bold indicates the top two results.
+Our investigation into the efficacy of QMamba for IQA tasks reveals two key insights through analysis of the KADID dataset (7 distortion types) and cross-dataset validation.
+As shown in Figure 3, t-SNE visualization demonstrates superior distortion discrimination of QMamba: 1) Tightly clustered features for each distortion type, 2) Clear separation between dissimilar artifacts, and 3) Minimal inter-class overlap compared to ViT’s partial merging and CNN/Swin architectures’ significant feature entanglement.
+This discriminative capability directly impacts practical performance.
+While QMamba demonstrates modest gains on simpler datasets such as LIVE and CSIQ, which contain only 4 to 5 distortion types, it achieves substantial improvements on more complex benchmarks like TID2013 and KADID, which include 24 to 25 distortion types.
+The architecture based on state-space modeling enables adaptive frequency processing through a selective scanning mechanism, dynamically emphasizing distortion-critical patterns while suppressing irrelevant features.
+This approach stands in contrast to convolutional networks with fixed receptive fields and Transformers that tend to over-mix local characteristics through global attention.
+Our findings indicate that QMamba is well-suited for new quality assessment scenarios involving complex and mixed distortions, such as those introduced by neural compression or generative models.
+In these cases, traditional architectures often fail to capture subtle or entangled artifacts, while the selective modeling in QMamba provides better adaptability and robustness.
+4.3 Efficient Transfer Learning for Mamba-Based IQA
+In the context of transferable IQA tasks, we conducted domain-specific training using synthetic datasets (LIVE, KADID), authentic datasets (LIVEC, KonIQ), and AIGC datasets (AIGC2023, AGIQA3K).
+After training in one domain, models were directly transferred and tested on datasets from the other two domains.
+We employed the StylePrompt technique, as illustrated in Figure 2, where the architecture was kept intact by freezing all model parameters and fine-tuning only the StylePrompt module, which involved approximately $4\%$ of the total parameter count.
+This approach achieved performance levels comparable to those obtained through full-parameter training.
+The outcomes, presented in Table 3, clearly demonstrate the effectiveness and efficiency of the proposed StylePrompt method for transferable IQA tasks.
+The value of ”Average” represents the mean performance across all domain transfers, with more detailed results provided in the Appendix.
+4.4 Ablation Study
+4.4.1 Different Scanning Method
+As shown in the task-specific results (Table 1) and universal evaluation results (Table 2), models adopting local scanning consistently outperform those based on cross-scanning strategies.
+This motivates our exploration of local scanning mechanisms, as employed in LQMamba.
+Although the average performance gap between QMamba and LQMamba may appear small, further analysis reveals that LQMamba performs better on most individual datasets.
+The marginal overall gain is primarily due to relatively lower improvements on simpler datasets such as LIVE and CSIQ, which contain fewer distortion types and less diverse content.
+In contrast, on more challenging datasets like TID2013 and KADID-10k, which feature a wide variety of fine-grained distortions, LQMamba shows clear advantages (e.g., SRCC on TID2013: 0.964 vs.
+0.949; on KADID: 0.941 vs.
+0.932).
+These results highlight the effectiveness of the local scanning design in complex, distortion-rich scenarios where precise modeling of local artifacts is critical.
+4.4.2 Different Model Scale
+In our empirical analysis, as documented in Table 1, we discern that while QMamba-Base exhibits exceptionally robust quality perception capabilities, QMamba-Small either matches or exceeds the performance of QMamba-Base across the majority of the datasets.
+Although QMamba-Tiny displays a modest decline in performance metrics, it still delivers results that are competitive with current SOTA methods, solely utilizing the capabilities of QMamba-Tiny.
+  Tuning Strategy Parameters PLCC_Avg. SRCC_Avg.
+  SSF 6.1M 0.750 0.735
+Crossattn_Prompt 12.17M 0.806 0.772
+Conv_Prompt 28.33M 0.883 0.856
+StylePrompt(ours) 3.83M 0.911 0.890
+Table 4: Ablation Study on Different Prompt Tuning Strategies
+4.4.3 Different Prompt Tuning Strategies
+To validate the effectiveness of the StylePrompt Generation (SPG) process, we conducted tests by directly learning a set of affine parameters $\gamma$ and $\beta$ to modulate the original features, rather than using prompts for learning, similar to the SSF(Lian et al., 2022).
+Additionally, to assess the effectiveness of the StylePrompt interaction method, we explored various interaction strategies during tuning, including convolutional prompt interaction and cross-attention prompt interaction.
+The results in Table 4 demonstrate the superior efficiency and performance of StylePrompt.
+In addition, we conducted several ablation studies related to the design of the prompts, including variations in the number and shape of the prompts.
+The detailed results of these experiments are provided in Appendix.
+4.5 Discussion and Future Work
+Looking ahead, the sequential modeling nature and efficiency of our SSM-based architecture make it a promising candidate for extension to video and audio quality assessment.
+Given the higher temporal complexity in VQA and the inherent sequential structure of audio signals, our method offers a unified and lightweight foundation for future research across visual and multimodal quality evaluation tasks.
+5 Conclusion
+In this paper, we introduced QMamba, a novel state space framework for image quality assessment that integrates task-specific evaluation, universal perception, and cross-domain transferability.
+Extensive experiments demonstrate that QMamba consistently outperforms established vision models, achieving significant improvements in accuracy while requiring substantially lower computational resources.
+The proposed StylePrompt mechanism enables robust cross-domain adaptation through lightweight dynamic feature recalibration, setting new benchmarks for IQA and demonstrating strong practical potential in real-world applications.
+Appendix A Appendix / supplemental material
+A.1 Detailed Description of the Datasets
+We evaluate the performance of the proposed method across eight widely recognized BIQA datasets, comprising both synthetic and authentic datasets.
+The synthetic datasets include LIVE(Sheikh et al., 2006), CSIQ(Larson & Chandler, 2010), TID2013(Ponomarenko et al., 2015), and KADID-10k(Lin et al., 2019).
+These datasets feature a small number of pristine images that are synthetically distorted using various techniques such as JPEG compression and Gaussian blurring.
+Specifically, LIVE contains 799 images with 5 types of distortion, CSIQ includes 866 images with 6 distortion types, TID2013 comprises 3000 images with 24 distortion types, and KADID-10k includes 10125 images with 25 distortion types.
+On the other hand, the authentic datasets include LIVEC(Ghadiyaram & Bovik, 2015), KonIQ-10k(Hosu et al., 2020), SPAQ(Fang et al., 2020), and FLIVE(Ying et al., 2020).
+LIVEC consists of 1,162 images with diverse authentic distortions captured by mobile devices.
+KonIQ-10k is composed of 10,073 images selected from YFCC-100M, covering a wide range of distortions such as brightness, colorfulness, contrast, noise, and sharpness.
+SPAQ contains 11,125 images collected by various smartphones, representing a large variety of scene categories.
+FLIVE, the largest in-the-wild IQA dataset to date, contains 39,810 real-world images with diverse content, sizes, and aspect ratios.
+In response to the rapid development of AI-generated content, we also employed two additional datasets: AIGCIQA2023(Wang et al., 2023a) and AGIQA3K(Li et al., 2023a).
+AIGCIQA2023 contains over 2000 images generated by six state-of-the-art text-to-image models, evaluated through a subjective experiment on quality, authenticity, and correspondence.
+AGIQA3K consists of 2,982 images from GAN, autoregression, and diffusion-based models, with annotations for perceptual quality and text-to-image alignment.
+A.2 Additional Experimental Results
+A.2.1 Task-specific IQA results on AIGC datasets
+We present the test results of various backbones on the AIGC datasets in Table 5, with bold indicating the best results.
+  Model AGIQA3K AIGCIQA2023
+  GFLOPS PLCC SRCC PLCC SRCC Average
+     ResNet-50(23.51M) 4.11G 0.901 0.840 0.795 0.797 0.833
+ResNet-101(42.50M) 7.83G 0.907 0.847 0.834 0.831 0.855
+ResNet-152(58.15M) 11.53G 0.901 0.832 0.841 0.834 0.852
+ViT-T(5.52M) 1.26G 0.865 0.787 0.760 0.766 0.795
+ViT-S(21.67M) 4.61G 0.891 0.819 0.842 0.822 0.844
+ViT-B(85.80M) 17.58G 0.897 0.830 0.853 0.835 0.854
+Swin-T(27.52M) 4.51G 0.906 0.847 0.867 0.844 0.866
+Swin-S(48.84M) 8.77G 0.908 0.849 0.875 0.857 0.872
+Swin-B(86.74M) 15.47G 0.909 0.852 0.886 0.863 0.878
+QMamba-T (27.99M) 4.47G 0.913 0.858 0.888 0.873 0.883
+QMamba-S (49.37M) 8.71G 0.912 0.858 0.889 0.875 0.884
+QMamba-B (87.53M) 15.35G 0.914 0.861 0.886 0.868 0.882
+LQMamba-T(29.87M) 4.44G 0.914 0.862 0.884 0.868 0.882
+LQMamba-S(52.91M) 8.66G 0.913 0.864 0.888 0.869 0.884
+LQMamba-B(93.79M) 15.30G 0.915 0.858 0.888 0.871 0.883
+Table 5: Results of AIGC for task-specific IQA on various backbones
+A.2.2 Universal IQA detailed data
+We provide the detailed results of Universal IQA in Table 6.
   Train LIVE & KADID & LIVEC & KonIQ & AGIQA3K & AIGCIQA2023
@@
      ResNet-50 4.11G 0.916 0.908 0.868 0.869 0.885 0.841 0.896 0.869 0.884 0.813 0.821 0.815 0.865
-ResNet-101 7.83G 0.927 0.921 0.870 0.863 0.885 0.848 0.906 0.878 0.897 0.841 0.843 0.833 0.876
-ResNet-152 11.53G 0.936 0.932 0.872 0.866 0.890 0.853 0.911 0.889 0.901 0.841 0.852 0.844 0.882
-ViT-T 1.26G 0.878 0.878 0.829 0.828 0.811 0.773 0.824 0.805 0.824 0.745 0.767 0.766 0.811
 ViT-S 4.61G 0.921 0.918 0.920 0.917 0.883 0.828 0.911 0.892 0.885 0.822 0.825 0.824 0.879
-ViT-B 17.58G 0.941 0.935 0.903 0.901 0.884 0.844 0.912 0.899 0.890 0.826 0.848 0.832 0.885
 Swin-T 4.51G 0.927 0.927 0.925 0.921 0.889 0.863 0.929 0.918 0.897 0.840 0.834 0.828 0.892
-Swin-S 8.77G 0.946 0.949 0.920 0.916 0.913 0.884 0.929 0.912 0.902 0.844 0.861 0.852 0.902
-Swin-B 15.47G 0.945 0.942 0.940 0.937 0.916 0.886 0.940 0.928 0.905 0.847 0.880 0.865 0.911
+DEIQT 4.68G 0.907 0.906 0.898 0.895 0.896 0.855 0.928 0.906 0.902 0.841 0.839 0.835 0.884
+LoDa 23.68G 0.895 0.902 0.907 0.900 0.872 0.842 0.900 0.875 0.874 0.809 0.809 0.801 0.866
 QMamba-T 4.47G 0.923 0.923 0.938 0.932 0.898 0.863 0.932 0.917 0.906 0.851 0.835 0.830 0.896
-QMamba-S 8.71G 0.938 0.941 0.944 0.938 0.905 0.875 0.937 0.923 0.909 0.855 0.853 0.841 0.905
-QMamba-B 15.35G 0.922 0.924 0.947 0.942 0.911 0.883 0.939 0.924 0.908 0.857 0.852 0.840 0.904
 LQMamba-T 4.44G 0.929 0.926 0.943 0.939 0.899 0.863 0.936 0.921 0.908 0.853 0.840 0.828 0.899
-LQMamba-S 8.66G 0.935 0.936 0.942 0.938 0.902 0.870 0.938 0.923 0.912 0.860 0.859 0.843 0.905
-LQMamba-B 15.30G 0.941 0.939 0.944 0.941 0.916 0.889 0.939 0.927 0.910 0.859 0.870 0.857 0.911
-Table 3: Performance comparison for transferable IQA.
-  Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
-  Test KADID LIVE AIGC2023 AGIQA3K KonIQ LIVEC AIGC2023 AGIQA3K KonIQ LIVEC KADID LIVE
-    Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
-    Without_tuning 0.535 0.501 0.726 0.744 0.789 0.784 0.726 0.744 0.553 0.549 0.595 0.550 0.600 0.612 0.675 0.649 0.636 0.587 0.667 0.637 0.460 0.425 0.815 0.838 0.642
-Lin_Probe(R) 0.606 0.581 0.783 0.811 0.824 0.804 0.769 0.695 0.871 0.848 0.785 0.751 0.803 0.788 0.771 0.723 0.818 0.792 0.790 0.761 0.570 0.540 0.799 0.826 0.755
+Table 6: Performance comparison for universal IQA.
+A.2.3 Transferable IQA results trained on the AIGC domain
+We provide the detailed results of Transferable IQA trained in the AIGC domain in Table 7.
+Train AIGC2023 & AGIQA3K (AIGC)
+Test KonIQ LIVEC KADID LIVE
+Fine-tuning Method PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC Average
+DEIQT 0.578 0.505 0.684 0.638 0.498 0.464 0.752 0.750 0.601
+LoDa 0.564 0.500 0.542 0.548 0.555 0.524 0.840 0.833 0.622
+Without_tuning 0.636 0.587 0.667 0.637 0.460 0.425 0.815 0.838 0.642
+Lin_Probe(R) 0.818 0.792 0.790 0.761 0.570 0.540 0.799 0.826 0.755
 Full_tuning (93.79M)
-0.936 0.930 0.941 0.940 0.885 0.863 0.910 0.851 0.943 0.928 0.899 0.873 0.880 0.859 0.910 0.856 0.943 0.927 0.910 0.878 0.937 0.932 0.938 0.933 0.908
+0.943 0.927 0.910 0.878 0.937 0.932 0.938 0.933 0.908
 StylePrompt (3.83M)
-0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
-StylePrompt & R 0.921 0.913 0.944 0.945 0.877 0.853 0.906 0.847 0.931 0.912 0.889 0.860 0.876 0.857 0.905 0.849 0.929 0.909 0.899 0.859 0.918 0.911 0.926 0.927 0.898
-4.3 Efficent transfer learning for Mamba-based IQA
-In the context of transferable IQA tasks, we engaged in domain-specific training utilizing synthetic datasets (LIVE, KADID), authentic datasets (LIVEC, KonIQ), and AIGC datasets (AIGC2023, AGIQA3K).
-Each dataset underwent subsequent evaluations in the other two domains.
-We employed the StylePrompt technique, whereby we maintained the architecture’s integrity during cross-domain transfer by freezing it entirely, and solely fine-tuned the StylePrompt, utilizing merely $4\%$ of the total parameter volume.
-This approach achieved results that were nearly equivalent to those obtained through full-parameter training.
-The outcomes, as delineated in Table 3, present a comparative analysis of various fine-tuning strategies, conclusively validating the efficacy and efficiency of our proposed StylePrompt method in facilitating transferable IQA tasks.
-4.4 Ablation Studies
-4.4.1 Different scanning
-Examining the outcomes from task-specific assessment in Table 1 and universal assessment in Table 2 , models equipped with local scanning have shown a definitive advantage over those utilizing cross-scanning techniques.
-This observation has led to an exploration of the effectiveness of local scanning methods.
-4.4.2 Different model size
-In our empirical analysis, as documented in Table 1 and Table 2, we discern that while QMamba-Base exhibits exceptionally robust quality perception capabilities, QMamba-Small either matches or exceeds the performance of QMamba-Base across the majority of the datasets.
-Although QMamba-Tiny displays a modest decline in performance metrics, it still delivers results that are competitive with current state-of-the-art (SOTA) methods, solely utilizing the capabilities of QMamba-Tiny.
-Table 4: Ablation study on different prompt tuning strategies
+0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
+StylePrompt & R 0.929 0.909 0.899 0.859 0.918 0.911 0.926 0.927 0.898
+Table 7: Performance comparison for transferable IQA (trained on AIGC).
+A.2.4 Detailed results of ablation studies
+  • Ablation study on different prompt tuning strategies.
+    We have provided the detailed experimental data for the ablation study on different prompt tuning strategies in Table 8.
+  • Ablation study results for different prompt designs.
+    We investigated the impact of the number of prompt components on performance and found that setting the number to six yields optimal results.
+Additionally, we explored whether varying the spatial dimensions and sizes of prompts would enhance performance.
+Our findings show that a spatial size of (1,1), focusing solely on the channel dimension, offers the best results.
+The outcomes of these ablation studies across multiple datasets are presented in Tables 9 and 10.
   Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
@@
 StylePrompt(ours)(3.83M) 0.920 0.912 0.949 0.948 0.877 0.854 0.908 0.854 0.932 0.913 0.888 0.866 0.880 0.865 0.906 0.852 0.929 0.909 0.902 0.871 0.918 0.913 0.926 0.922 0.901
-4.4.3 Different prompt tuning strategies
-To validate the effectiveness of the StylePrompt Generation (SPG) process, we tested the direct learning of a set of affine parameters, $\gamma$ and $\beta$, to modulate the original features.
-The results in Table 4 demonstrate the outcomes of transfer learning using only the learned affine parameters, substantiating the effectiveness of our StylePrompt Generation phase.
-Furthermore, to verify the efficacy of different prompt interaction methods within the StylePrompt, we evaluated various prompt interaction strategies with the original features.
-These included convolutional prompt interaction and cross-attention prompt interaction.
-Experimental results indicated that the interaction method employed by the StylePrompt with the original features achieves the best transfer results with minimal parameter usage.
-This further validates the effectiveness of our StylePrompt Injection (SPI) phase.
-Table 5: Ablation study on the number of prompt
-  Num of Prompt PLCC_Average SRCC_Average Average
-  N=1 0.905 0.881 0.8928
-N=2 0.905 0.882 0.8936
-N=4 0.907 0.884 0.8951
-N=6 0.911 0.890 0.9006
-N=8 0.905 0.883 0.8939
-N=10 0.906 0.883 0.8942
-Table 6: Ablation study on the shape of prompts
-  Prompt_shape(H,W) PLCC_Average SRCC_Average Average
-  (1,1) 0.911 0.890 0.901
-(7,7) 0.909 0.886 0.897
-(14,14) 0.909 0.886 0.897
-(28,28) 0.907 0.883 0.895
-layer_HW 0.908 0.884 0.896
-4.4.4 Different Prompt Designs
-We investigated the impact of the number of prompt components on performance.
-Our findings revealed that setting the number of prompts to six yields optimal performance.
-Furthermore, we explored whether prompts with different spatial dimensions and sizes contribute to enhanced capabilities.
-Ultimately, we found that a spatial size of (1,1), which solely consists of the channel dimension, yields superior performance.
-We report the average results of these two ablation studies across multiple datasets in Tables 5 and Table 6.
-More detailed data will be provided in the appendix Table 8 and Table 9.
-5 Conclusion
-In this paper, we introduced Q-Mamba, a novel exploration of the Mamba model in the IQA field, which aims to enhance image quality assessment (IQA) from three perspectives: task-specific, universal, and transferable IQA.
-Our extensive experiments have demonstrated that Q-Mamba has great potential to outperform existing foundational models in IQA, including Swin Transformer, ViT, and CNNs, in both perceptual accuracy and computational efficiency.
-To improve the perception transferability of Q-Mamba, we developed StylePrompt, a powerful tuning mechanism, to enable the effective perception knowledge transfer of Q-Mamba to various downstream IQA tasks even with large domain shifts, by exploiting only extra $4\%$ of whole parameters to adjust the statistics of perception features of Q-Mamba.
-We believe these findings can prompt the further development of IQA and pave the way for future research of Mamba-based quality assessment.
-Appendix A Appendix / supplemental material
-A.1 Additional experimental results
-A.1.1 Task-specific results on AIGCIQA
-We present the test results of various backbones on the AIGC datasets in Table 7, with bold indicating the best results.
-Table 7: Results of AIGC on various backbones
-  Model AGIQA3K AIGCIQA2023
-  FLOPs PLCC SRCC PLCC SRCC PLCC_Average SRCC_Average Average
-     ResNet-50(23.51M) 4.11G 0.901 0.840 0.795 0.797 0.848 0.819 0.833
-ResNet-101(42.50M) 7.83G 0.907 0.847 0.834 0.831 0.871 0.839 0.855
-ResNet-152(58.15M) 11.53G 0.901 0.832 0.841 0.834 0.871 0.833 0.852
-ViT-T(5.52M) 1.26G 0.865 0.787 0.760 0.766 0.813 0.777 0.795
-ViT-S(21.67M) 4.61G 0.891 0.819 0.842 0.822 0.867 0.821 0.844
-ViT-B(85.80M) 17.58G 0.897 0.830 0.853 0.835 0.875 0.833 0.854
-Swin-T(27.52M) 4.51G 0.906 0.847 0.867 0.844 0.887 0.846 0.866
-Swin-S(48.84M) 8.77G 0.908 0.849 0.875 0.857 0.892 0.853 0.872
-Swin-B(86.74M) 15.47G 0.909 0.852 0.886 0.863 0.898 0.858 0.878
-QMamba-T (27.99M) 4.47G 0.913 0.858 0.888 0.873 0.901 0.866 0.883
-QMamba-S (49.37M) 8.71G 0.912 0.858 0.889 0.875 0.901 0.867 0.884
-QMamba-B (87.53M) 15.35G 0.914 0.861 0.886 0.868 0.900 0.865 0.882
-LQMamba-T(29.87M) 4.44G 0.914 0.862 0.884 0.868 0.899 0.865 0.882
-LQMamba-S(52.91M) 8.66G 0.913 0.864 0.888 0.869 0.901 0.867 0.884
-LQMamba-B(93.79M) 15.30G 0.915 0.858 0.888 0.871 0.902 0.865 0.883
-A.1.2 Detailed Results of Ablation Studies
-We report the detailed data of our experiments on Prompt design in Table 8 and Table 9.
-Table 8: Ablation study results for different numbers of prompts (N).
+Table 8: Ablation study on different prompt tuning strategies
   Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
@@
 N=10 0.919 0.913 0.930 0.936 0.877 0.856 0.901 0.840 0.929 0.910 0.880 0.849 0.877 0.853 0.900 0.840 0.925 0.906 0.889 0.851 0.918 0.914 0.923 0.924 0.8942
-Table 9: Ablation study for different prompt shape
+Table 9: Ablation study results for different numbers of prompts (N).
   Train KonIQ & LIVEC KADID & LIVE AIGC2023 & AGIQA3K
@@
 layer_HW 0.921 0.914 0.948 0.948 0.880 0.854 0.905 0.844 0.929 0.909 0.876 0.850 0.879 0.859 0.903 0.844 0.928 0.906 0.891 0.858 0.913 0.908 0.920 0.919 0.896
-A.2 Limitations
+Table 10: Ablation study for different prompt shape
+A.3 Limitations
 Q-Mamba heavily relies on pre-trained weights from ImageNet-1K, which may limit its applicability to domains with significantly different data distributions.
 Future work could explore pre-training on more diverse datasets to improve generalization capabilities.
-Despite the efficiency improvements brought by the StylePrompt, the computational demands for training and deploying Q-Mamba on very large-scale datasets or in real-time applications might still be substantial.
+Despite the efficiency improvements brought by the StylePrompt, the computational demands for training and deploying QMamba on very large-scale datasets or in real-time applications might still be substantial.
 Investigating methods to further reduce computational complexity without sacrificing performance could be beneficial.
 The current study focuses on specific types of distortions, and there may be other types of distortions that have not been sufficiently explored.
-Expanding the evaluation to cover a broader range of distortions could provide a more comprehensive validation of Q-Mamba’s robustness.
+Expanding the evaluation to cover a broader range of distortions could provide a more comprehensive validation of QMamba’s robustness.
 Given that synthetic datasets might not fully capture the complexity and variability of authentic data, there is a risk of overfitting to these synthetic examples.
 Further evaluations on more diverse and extensive authentic datasets would help ensure the model’s robustness and practical applicability.
-While Q-Mamba shows promising results in cross-domain transferability, its effectiveness across vastly different domains (e.g., medical imaging versus natural images) has not been thoroughly validated.
-Further studies are needed to test and possibly adapt Q-Mamba for such diverse applications.
-A.3 Broader Impacts
-Improving image quality assessment (IQA) through Q-Mamba allows users across various platforms to enjoy higher quality images, enhancing user experience in social media, streaming services, and other digital platforms, thus leading to greater satisfaction and engagement.
-The enhancement of image quality can significantly benefit individuals with visual impairments, as better image clarity and quality make content more accessible, thereby improving the inclusivity of digital media.
-Additionally, the computational efficiency of Q-Mamba can result in more efficient image processing workflows, reducing the time and energy required for image enhancement, compression, and transmission.
-This can lead to cost savings and environmental benefits due to lower energy consumption.
-Moreover, the transferability of Q-Mamba across different domains can improve image quality in fields such as medical imaging, satellite imagery, and autonomous driving, thereby enhancing the accuracy and reliability of critical applications.
-While Q-Mamba presents these positive societal impacts, it is important to acknowledge and mitigate potential negative effects.
-Enhancing image quality and clarity may inadvertently lead to privacy concerns, as higher quality images can capture more detailed information about individuals and environments.
-However, our work adheres to ethical guidelines and uses publicly available, open-source datasets, ensuring compliance with various constraints and regulations.
-By focusing on responsible and ethical use, we aim to maximize the benefits of Q-Mamba while minimizing any adverse societal impacts.
+While QMamba shows promising results in cross-domain transferability, its effectiveness across vastly different domains (e.g., medical imaging versus natural images) has not been thoroughly validated.
+Further studies are needed to test and possibly adapt QMamba for such diverse applications.